### v1

# Error-Correcting Codes for Two Bursts of $t_{1}$-Deletion-$t_{2}$-Insertion with Low Computational Complexity

Abstract
Burst errors involving simultaneous insertions, deletions, and substitutions occur in practical scenarios, including DNA data storage and document synchronization, motivating developments of channel codes that can correct such errors. In this paper, we address the problem of constructing error-correcting codes (ECCs) capable of handling multiple bursts of $t_{1}$-deletion-$t_{2}$-insertion ($(t_{1},t_{2})$-DI) errors, where each burst consists of $t_{1}$ deletions followed by $t_{2}$ insertions in a binary sequence. We make three key contributions: Firstly, we establish the fundamental equivalence of (1) two bursts of $(t_{1},t_{2})$-DI ECCs, (2) two bursts of $(t_{2},t_{1})$-DI ECCs, and (3) one burst each of $(t_{1},t_{2})$-DI and $(t_{2},t_{1})$-DI ECCs. Then, we derive lower and upper bounds on the code size of two bursts of $(t_{1},t_{2})$-DI ECCs, which can naturally be extended to the case of multiple bursts. Finally, we present constructions of two bursts of $(t_{1},t_{2})$-DI ECCs. Compared to the codes obtained by the syndrome compression technique, the resulting codes achieve significantly lower computational complexity.
I Introduction
Channels with insertions and deletions, dubbed as synchronization error channels, model certain communication systems, including magnetic recording channels [12], document synchronization applications [5], and DNA data storage systems [8]. Over the past half century, there have been many developments addressing different aspects of such systems. For instance, different types of channel codes have been developed for correcting deletion and insertion errors [11, 2, 9, 16]. Despite all the developments over the years on the very basic models for deletion/insertion errors, there are many other scenarios that have not been addressed. Of particular interest is the presence of insertion, deletion, and substitution errors occurring together. With this motivation, there have been recent works on developing error-correcting codes (ECCs) capable of effectively addressing more general channels with synchronization errors.
Despite many recent works, the development of ECCs for synchronization errors is a fundamental challenge. Going back to the foundations, the seminal work by Levenshtein [11] constructed a single-edit (insertion/deletion/substitution) correcting code requiring only $\log n+1$ bits of redundancy, using the well-known Varshamov-Tenengolts (VT) syndrome [33]. [^1] In this paper, unless stated otherwise, all logarithmic operations are base two. While there were only sporadic results for a long time, recent breakthroughs by Brakensiek et al. [3] on $t$-deletion codes have spurred substantial progress in this field [6, 25, 7, 23, 21, 22, 17, 30, 4, 28, 13, 29, 19] over the last decade. For the specific case of two deletions, the state-of-the-art result was achieved by Guruswami et al. [7] for binary codes with $4\log n+o(\log n)$ bits of redundancy. This result was extended to non-binary alphabets by Song et al. [28] and Liu et al. [13]. They introduced codes with $5\log n+o(\log n)$ bits of redundancy under certain parameter constraints. For general $t$-deletions with $t\geq 2$, Sima et al. [21, 23] developed binary codes with $4t\log n+o(\log n)$ bits of redundancy using the syndrome compression technique [24]. Subsequent work by Song et al. [29] improved this result to $(4t-1)\log n+o(\log n)$ through pre-coding methods. Regarding the $q$-ary case, Sima et al. [22] initially achieved at most $30t\log n+o(\log n)$ bits of redundancy for $q$-ary $t$-deletion codes, and Schaller et al. [19] recently improved this result to $t\log q+(3t-1)\log n+o(\log n)$ bits of redundancy for sufficiently large $n$ and $q>n^{2+\epsilon}$, where $\epsilon>0$.
In some practical scenarios, data corruption may occur as bursts of insertions, deletions, and substitutions. Motivated by this, significant research has focused on designing ECCs capable of handling such errors as well. Several works have investigated code constructions for correcting a single burst of $t$-deletions or up to $t$-deletions, including [20, 10, 28, 27, 34]. Among these, the recent coding scheme proposed by Song et al. [27] achieves a redundancy of $\log n+8\log\log n+o(\log\log n)$ bits for correcting a single burst of at most $t$-deletions. As a further extension in this direction, Sun et al. [31] proposed ECCs with $\log n+o(\log n)$ bits of redundancy to correct a single burst of edits in $q$-ary sequences ($q\geq 2$). Beyond the single-burst setting, Sima et al. [24] constructed ECCs capable of correcting $m$ bursts of at most $t$-deletions with a redundancy of $4m\log n+o(\log n)$ bits by employing the syndrome compression technique. After that, focusing on the case of two bursts of exactly $t$-deletions, Ye et al. [35] presented explicit ECC constructions with $5\log n+o(\log n)$ bits of redundancy.
In real-world applications, multiple error types may occur simultaneously. Early work by Smagloy et al. [26] addressed this challenge by constructing binary ECCs correcting a single deletion and a single substitution. Later, Song et al. [29] generalized this result to multiple deletions and multiple substitutions under both binary and non-binary settings. A notable contribution by Schoeny et al. [20] introduced burst ECCs capable of handling both deletions and insertions simultaneously, specifically constructing binary codes for one burst of 22-deletions-11-insertion (denoted by $(2,1)$-DI) errors. This result was later extended by Lu et al. to one burst of $(t_{1},1)$-DI ECCs [14] and subsequently to $(t_{1},t_{2})$-DI ECCs with $t_{1}\geq 2t_{2}$ [15], thereby establishing a lower bound of $\log(n-t_{1}+2)+t_{1}-1$ bits of redundancy for such codes. Most recently, Sun et al. [32] provided explicit constructions of $q$-ary one burst of $(t_{1},t_{2})$-DI ECCs with $\log n+O(1)$ bits of redundancy, applicable for any non-negative integers $t_{1},t_{2}$ and $q\geq 2$.
Building upon the developments for correcting different types of burst errors, in this paper, we first establish the equivalence of two bursts of $(t_{1},t_{2})$-DI ECCs, two bursts of $(t_{2},t_{1})$-DI ECCs, and one burst of $(t_{1},t_{2})$-DI one burst of $(t_{2},t_{1})$-DI ECCs. We then derive both lower and upper bounds on the cardinality of two bursts of $(t_{1},t_{2})$-DI ECCs. Based on these results, we further propose novel constructions of binary ECCs capable of correcting two bursts of $(t_{1},t_{2})$-DI with $t_{1},t_{2}$ being two positive integers. The resulting codes achieve substantially reduced computational complexity compared to those obtained by the syndrome compression technique.
The remainder of this paper is organized as follows. Section II introduces some necessary notation, definitions and preliminary observations. In Section III, we first establish the equivalence among two bursts of $(t_{1},t_{2})$-DI ECCs, two bursts of $(t_{2},t_{1})$-DI ECCs, and one burst of $(t_{1},t_{2})$-DI one burst of $(t_{2},t_{1})$-DI ECCs, and then analyze the lower and upper bounds on the cardinality of two bursts of $(t_{1},t_{2})$-DI ECCs. The constructions of two bursts of $(t_{1},t_{2})$-DI ECCs are developed in Sections IV and V. Finally, Section VI concludes the paper.
II Preliminaries
In this section, we provide some necessary definitions and previously known results, which will be applied in the sequel to construct two bursts of $(t_{1},t_{2})$-DI ECCs. For convenience, we first summarize the notation used throughout the paper.
For two non-negative integers $a$ and $b$ with $a<b$, define two ordered sets $\{a,a+1,\dots,b-1\}$ and $\{a,a+1,\dots,b\}$ by $[a,b)$ and $[a,b]$, respectively. When $a=1$, we abbreviate $[1,b]$ as $[b]$. The cardinality of a set $\cal{A}$ is denoted by $|\mathcal{A}|$. For an $m\times n$ matrix $A$, denote the $i$-th row by $A[i],i\in[m]$. For any positive integer $q\geq 2$, denote $\Sigma_{q}=\{0,1,\dots,q-1\}$. Let $|\mathbf{x}|=n$ be the length of $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})\in\Sigma_{2}^{n}$. For an index set $\mathcal{I}=\{i_{1},i_{2},\dots,i_{j}\}\subseteq[n]$, $\mathbf{x}_{\mathcal{I}}=(x_{i_{1}},x_{i_{2}},\dots,x_{i_{j}})$ represents the projection of $\mathbf{x}$ onto the indices of $\mathcal{I}$. In this case, $\mathbf{x}_{\mathcal{I}}$ is referred to as a substring of $\mathbf{x}$ when $i_{j^{\prime}-1}=i_{j^{\prime}}-1$ for all $j^{\prime}\in[2,j]$. If $x_{i}=x_{i+1}$ for any $i\in[i_{1},i_{j}-1]$, we call the substring $(x_{i_{1}},x_{i_{2}},\dots,x_{i_{j}})$ a run of $\mathbf{x}$, for which $x_{i_{1}-1}\neq x_{i_{1}},x_{i_{j}}\neq x_{i_{j}+1}$ and we default $x_{1}\neq x_{0},x_{n}\neq x_{n+1}$. If $x_{i}=x_{i+2}$ for any $i\in[i_{1},i_{j}-2]$, it forms an alternating substring in $\mathbf{x}$.
Definition 1.
For a binary sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})\in\Sigma_{2}^{n}$ and two non-negative integers $t_{1},t_{2}$, if the substring $(x_{i},x_{i+1},\dots,x_{i+t_{1}-1})$ is deleted and then $(b_{1},b_{2},\dots,b_{t_{2}})$ is inserted at the $i$-th coordinate of $\mathbf{x}$, where $b_{1}\neq x_{i},b_{t_{2}}\neq x_{i+t_{1}-1}$ for $i\in[n-t_{1}+1]$, we say that a one-burst of $(t_{1},t_{2})$-DI occurred at the $i$-th component of $\mathbf{x}$.
Definition 2.
For a binary sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})\in\Sigma_{2}^{n}$ and two non-negative integers $t_{1},t_{2}$, if the substring $(x_{i},x_{i+1},\dots,x_{i+t_{1}-1})$ is deleted and then $(x_{i+t_{1}},$ $x_{i+t_{1}+1},\dots,x_{i+t_{1}+t_{2}-1})$ is substituted by $(b_{1},b_{2},\dots,b_{t_{2}})$, where $b_{1}\neq x_{i}$ for $i\in[n-t_{1}+1]$, we say that a one-burst of $t_{1}$-deletions-$t_{2}$-substitutions ($(t_{1},t_{2})$-DS) occurred at the $i$-th component of $\mathbf{x}$.
Denote by $\mathcal{B}_{m,(t_{1},t_{2})}^{DI}(\mathbf{x})$ and $\mathcal{B}_{m,(t_{1},t_{2})}^{DS}(\mathbf{x})$ the sets of sequences obtained from $\mathbf{x}$ by $m$ bursts of $(t_{1},t_{2})$-DI and $m$ bursts of $(t_{1},t_{2})$-DS, respectively. We simply have $\mathcal{B}_{(t_{1},t_{2})}^{DI}(\mathbf{x})$ and $\mathcal{B}_{(t_{1},t_{2})}^{DS}(\mathbf{x})$ when $m=1$.
Thus, if $\mathbf{x}^{\prime}\in\mathcal{B}_{(t_{1},t_{2})}^{DI}(\mathbf{x})$, for $i\in[n-t_{1}+1]$, we can write
$\mathbf{x}^{\prime}=(x_{1},x_{2},\dots,x_{i-1},b_{1},b_{2},\dots,b_{t_{2}},x_{i+t_{1}},\dots,x_{n}).$
If $\mathbf{x}^{\prime\prime}\in\mathcal{B}_{(t_{1},t_{2})}^{DS}(\mathbf{x})$, we can write
$\mathbf{x}^{\prime\prime}=(x_{1},x_{2},\dots,x_{i-1},b_{1},b_{2},\dots,b_{t_{2}},x_{i+t_{1}+t_{2}},\dots,x_{n}).$
We remark that for one burst of $(t_{1},t_{2})$-DS errors, it is not necessary to set $x_{i+t_{1}+m-1}\neq b_{m}$ for any $m\in[t_{2}]$. This means that the $t_{2}$-substitution errors do not have to be contiguous, i.e., there are exactly $t_{1}$-deletions and up to $t_{2}$-substitutions in $\mathbf{x}^{\prime\prime}$.
Next, we formally define ECCs capable of correcting multiple bursts of $(t_{1},t_{2})$-DI. It is crucial to emphasize that the model developed throughout this work assumes that the presence of $m$ non-overlapping burst errors in the sequence.
Definition 3.
A subset of $\Sigma_{2}^{n}$, denoted by $\mathcal{C}$, is referred to as an $m$ bursts of $(t_{1},t_{2})$-DI ECC if it is capable of correcting $m$ bursts of $(t_{1},t_{2})$-DI. That is, for any two distinct sequences $\mathbf{x},\mathbf{y}\in\mathcal{C}$, we have
$\mathcal{B}_{m,(t_{1},t_{2})}^{DI}(\mathbf{x})\cap\mathcal{B}_{m,(t_{1},t_{2})}^{DI}(\mathbf{y})=\varnothing.$
In what follows, we define $d$-regular sequences, which serve as a critical component throughout the paper.
Definition 4.
A sequence $\mathbf{x}\in\Sigma_{2}^{n}$ is referred to as a $d$-regular sequence if each substring of $\mathbf{x}$ with length at least $d\log n$ contains both $00$ and $11$.
In Definition 4, the parameter $d$ is a constant that can be chosen appropriately. Throughout this paper, we fix $d=7$. The following lemma presents an encoding method for constructing $d$-regular sequences with only one bit of redundancy.
Lemma 1 (Lemma 11, [7]).
For an integer $M\geq 2^{n-1}$, there exists a function $\mathrm{RegEnc}:[M]\rightarrow\Sigma_{2}^{n}$, which can be computed in near-linear time with a polynomial size lookup table, such that $\mathrm{RegEnc}(\cdot)$ is a $d$-regular sequence.
Observation 1.
If $\mathbf{x}\in\Sigma_{2}^{n}$ is a $d$-regular sequence, then the length of each run and alternating substring in $\mathbf{x}$ is at most $d\log n$ since it contains both $00$ and $11$.
III Equivalence of Different Burst ECCs and Bounds on their Cardinality
In this section, we first establish that ECCs capable of correcting two bursts of $(t_{1},t_{2})$-DI are equivalent to those correcting two bursts of $(t_{2},t_{1})$-DI, as well as to ECCs correcting one burst of $(t_{1},t_{2})$-DI together with one burst of $(t_{2},t_{1})$-DI. Subsequently, we derive lower and upper bounds on the cardinality of two bursts of $(t_{1},t_{2})$-DI ECCs.
III-A The Equivalence of Different Burst ECCs
We characterize the equivalence of different burst ECCs in the following two Theorems.
Theorem 1.
Let $t_{1}\geq t_{2}$ be two non-negative integers. A code $\mathcal{C}$ can correct two bursts of $(t_{1},t_{2})$-DI if and only if it can correct two bursts of $(t_{2},t_{1})$-DI.
Proof.
We show the ‘if’ direction by contradiction, and note that the ‘only if’ part can be proved similarly. An ECC $\mathcal{C}$ can correct two bursts of $(t_{1},t_{2})$-DI, this means that, for any $\mathbf{x},\mathbf{y}\in\mathcal{C}$, we have
$\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})\cap\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{y})=\varnothing.$
Assume that $\mathcal{C}$ cannot correct two bursts of $(t_{2},t_{1})$-DI. Then, there exist two different sequences $\mathbf{x},\mathbf{y}\in\mathcal{C}$ such that
$\mathcal{B}_{2,(t_{2},t_{1})}^{DI}(\mathbf{x})\cap\mathcal{B}_{2,(t_{2},t_{1})}^{DI}(\mathbf{y})\neq\varnothing.$
That is to say, we can find a sequence $\mathbf{s}\in\Sigma_{2}^{n-2t_{2}+2t_{1}}$ such that
$\mathbf{s}\in\mathcal{B}_{2,(t_{2},t_{1})}^{DI}(\mathbf{x})\cap\mathcal{B}_{2,(t_{2},t_{1})}^{DI}(\mathbf{y}).$
Suppose that $\mathbf{s}$ is obtained from $\mathbf{x}$ by deleting $(x_{i_{1}},x_{i_{1}+1},$ $\dots,x_{i_{1}+t_{2}-1})$ and $(x_{i_{2}},x_{i_{2}+1},\dots,x_{i_{2}+t_{2}-1})$, where $i_{2}\geq i_{1}+t_{2}$ ensures the separation between deletion positions, and subsequently inserting $(b_{1},b_{2},\dots,b_{t_{1}})$ and $(b^{\prime}_{1},b^{\prime}_{2},\dots,b^{\prime}_{t_{1}})$ at the $i_{1}$-th and $i_{2}$-th components of the original sequence $\mathbf{x}$, then we get
$\mathbf{s}=$ $x_{1},\dots,x_{i_{1}-1},b_{1},b_{2},\dots,b_{t_{1}},x_{i_{1}+t_{2}},\dots,x_{i_{2}-1},$
$b^{\prime}_{1},b^{\prime}_{2},\dots,b^{\prime}_{t_{1}},x_{i_{2}+t_{2}},\dots,x_{n}.$ (1)
Meanwhile, if $\mathbf{s}$ is obtained from $\mathbf{y}$ by deleting $(y_{j_{1}},y_{j_{1}+1},$ $\dots,y_{j_{1}+t_{2}-1})$ and $(y_{j_{2}},y_{j_{2}+1},\dots,y_{j_{2}+t_{2}-1})$, where $j_{2}\geq j_{1}+t_{2}$, and subsequently inserting $(c_{1},c_{2},\dots,c_{t_{1}})$ and $(c^{\prime}_{1},c^{\prime}_{2},\dots,c^{\prime}_{t_{1}})$ at the $j_{1}$-th and $j_{2}$-th components of the original sequence $\mathbf{y}$, then we get
$\mathbf{s}=$ $y_{1},\dots,y_{j_{1}-1},c_{1},c_{2},\dots,c_{t_{1}},y_{j_{1}+t_{2}},\dots y_{j_{2}-1},$
$c^{\prime}_{1},c^{\prime}_{2},\dots,c^{\prime}_{t_{1}},y_{j_{2}+t_{2}},\dots,y_{n}.$ (2)
In the following, we consider two cases according to whether the positions of two bursts in $\mathbf{x}$ and $\mathbf{y}$ have an intersection. To clarify, we show the two cases in Figure 1.
Figure 1: Cases i) and ii): two bursts of $(t_{1},t_{2})$-DI in $\mathbf{x}$ (blue) and $\mathbf{y}$ (red).
Case i) $i_{1}+t_{2}\leq i_{2}\leq j_{1}-2t_{1}+t_{2}\leq j_{2}-2t_{1}$. By comparing (III-A) and (III-A), we obtain
$(x_{1},\dots,x_{i_{1}-1})=(y_{1},\dots,y_{i_{1}-1}),$
$(b_{1},\dots,b_{t_{1}})=(y_{i_{1}},\dots,y_{i_{1}+t_{1}-1}),$
$(x_{i_{1}+t_{2}},\dots,x_{i_{2}-1})=(y_{i_{1}+t_{1}},\dots,y_{i_{2}+t_{1}-t_{2}-1}),$
$(b^{\prime}_{1},\dots,b^{\prime}_{t_{1}})=(y_{i_{2}+t_{1}-t_{2}},\dots,y_{i_{2}+2t_{1}-t_{2}-1}),$ (3)
$(x_{i_{2}+t_{2}},\dots,x_{j_{1}+2t_{2}-2t_{1}-1})=(y_{i_{2}+2t_{1}-t_{2}},\dots,y_{j_{1}-1}),$
$(x_{j_{1}+2t_{2}-2t_{1}},\dots,x_{j_{1}+2t_{2}-t_{1}-1})=(c_{1},\dots,c_{t_{1}}),$
$(x_{j_{1}+2t_{2}-t_{1}},\dots,x_{j_{2}+t_{2}-t_{1}-1})=(y_{j_{1}+t_{2}},\dots,y_{j_{2}-1}),$
$(x_{j_{2}+t_{2}-t_{1}},\dots,x_{j_{2}+t_{2}-1})=(c^{\prime}_{1},\dots,c^{\prime}_{t_{1}}),$
$(x_{j_{2}+t_{2}},\dots,x_{n})=(y_{j_{2}+t_{2}},\dots,y_{n}).$
Therefore, by deleting $(x_{j_{1}+2t_{2}-2t_{1}},\dots,x_{j_{1}+2t_{2}-t_{1}-1})$ and $(x_{j_{2}+t_{2}-t_{1}},\dots,x_{j_{2}+t_{2}-1})$ from $\mathbf{x}$, then inserting $(y_{j_{1}},\dots,$ $y_{j_{1}+t_{2}-1})$ and $(y_{j_{2}},\dots,y_{j_{2}+t_{2}-1})$ at the $j_{1}+2t_{2}-2t_{1}$-th and $j_{2}+t_{2}-t_{1}$-th components of $\mathbf{x}$, respectively, we can obtain
$\mathbf{x}^{\prime}=(x_{1},\dots,x_{j_{1}+2t_{2}-2t_{1}-1},y_{j_{1}},\dots,y_{j_{1}+t_{2}-1},x_{j_{1}+2t_{2}-t_{1}},$
$\dots,x_{j_{2}+t_{2}-t_{1}-1},y_{j_{2}},\dots,y_{j_{2}+t_{2}-1},x_{j_{2}+t_{2}},\dots,x_{n}).$
By deleting $(y_{i_{1}},\dots,y_{i_{1}+t_{1}-1})$ and $(y_{i_{2}+t_{1}-t_{2}},\dots,$ $y_{i_{2}+2t_{1}-t_{2}-1})$ from $\mathbf{y}$, then inserting $(x_{i_{1}},\dots,x_{i_{1}+t_{2}-1})$ and $(x_{i_{2}},\dots,x_{i_{2}+t_{2}-1})$ at the $i_{1}$-th and $i_{2}+t_{1}-t_{2}$-th components of $\mathbf{y}$, respectively, we can obtain
$\mathbf{y}^{\prime}=$ $(y_{1},\dots,y_{i_{1}-1},x_{i_{1}},\dots,x_{i_{1}+t_{2}-1},y_{i_{1}+t_{1}},\dots,$
$y_{i_{2}+t_{1}-t_{2}-1},x_{i_{2}},\dots,x_{i_{2}+t_{2}-1},y_{i_{2}+2t_{1}-t_{2}},\dots,y_{n}).$
With the equalities in (III-A), $\mathbf{x}^{\prime}=\mathbf{y}^{\prime}$ follows. This implies that
$\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})\cap\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{y})\neq\varnothing,$
which contradicts with the fact that $\mathcal{C}$ is an ECC for correcting two bursts of $(t_{1},t_{2})$-DI.
Case ii) $i_{1}+t_{2}\leq i_{2}\leq j_{1}\leq j_{2}-t_{2}~(j_{1}<i_{2}+2t_{1}-t_{2})$. By comparing (III-A) and (III-A), we obtain
[…]
Therefore, by deleting $(x_{j_{1}+t_{2}-2t_{1}},\dots,x_{j_{1}+t_{2}-t_{1}-1})$ and $(x_{j_{2}+t_{2}-t_{1}},\dots,x_{j_{2}+t_{2}-1})$ from $\mathbf{x}$, then inserting $(y_{j_{1}-t_{1}},\dots,y_{j_{1}+t_{2}-t_{1}-1})$ and $(y_{j_{2}},\dots,y_{j_{2}+t_{2}-1})$ at the $j_{1}+t_{2}-2t_{1}$-th and $j_{2}+t_{2}-t_{1}$-th components of $\mathbf{x}$, respectively, we can obtain
$\mathbf{x}^{\prime\prime}=($ $x_{1},\dots,x_{j_{1}+t_{2}-2t_{1}-1},y_{j_{1}-t_{1}},\dots,y_{j_{1}+t_{2}-t_{1}-1},$
$x_{j_{1}+t_{2}-t_{1}},\dots,x_{j_{2}+t_{2}-t_{1}-1},y_{j_{2}},\dots,y_{j_{2}+t_{2}-1},$
$x_{j_{2}+t_{2}},\dots,x_{n}).$
By deleting $(y_{i_{1}},\dots,y_{i_{1}+t_{1}-1})$ and $(y_{j_{1}+t_{2}-t_{1}},\dots,y_{j_{1}+t_{2}-1})$ from $\mathbf{y}$, then inserting $(x_{i_{1}},\dots,x_{i_{1}+t_{2}-1})$ and $(x_{j_{1}+t_{2}-t_{1}},$ $\dots,x_{j_{1}+2t_{2}-t_{1}-1})$ at the $i_{1}$-th and $j_{1}+t_{2}-t_{1}$-th components of $\mathbf{y}$, respectively, we can obtain
$\mathbf{y}^{\prime\prime}=($ $y_{1},\dots,y_{i_{1}-1},x_{i_{1}},\dots,x_{i_{1}+t_{2}-1},y_{i_{1}+t_{1}},\dots,$
$y_{j_{1}+t_{2}-t_{1}-1},x_{j_{1}+t_{2}-t_{1}},\dots,x_{j_{1}+2t_{2}-t_{1}-1},y_{j_{1}+t_{2}},$
$\dots,y_{n}).$
Since $i_{2}\leq j_{1}<i_{2}+2t_{1}-t_{2}$, we know $(x_{i_{1}+t_{2}},\dots,$ $x_{j_{1}+t_{2}-2t_{1}})=(y_{i_{1}+t_{1}},\dots,y_{j_{1}-t_{1}})$. With the equalities in (10), $\mathbf{x}^{\prime\prime}=\mathbf{y}^{\prime\prime}$ holds. This implies that
$\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})\cap\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{y})\neq\varnothing,$
which contradicts with the fact that $\mathcal{C}$ is an ECC for correcting two bursts of $(t_{1},t_{2})$-DI.
We note that additional cases may arise depending on the relative positions of two bursts in $\mathbf{x}$ and $\mathbf{y}$, for example, $i_{1}+t_{1}\leq j_{1}\leq i_{2}-t_{1}\leq j_{2}-2t_{1}$, or $i_{1}+t_{1}\leq j_{1}\leq j_{2}-t_{2}\leq i_{2}-t_{1}-t_{2}$, and so on. These cases can be proved analogously to Cases i) and ii). By comparing (III-A) and (III-A) under these configurations, we can establish the corresponding equivalence among the symbols, similar to those in (III-A) and (10). Based upon this, by first deleting and then inserting two substring into $\mathbf{x}$ and $\mathbf{y}$, it is evident that
$\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})\cap\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{y})\neq\varnothing,$
which leads to a contradiction. This completes the proof. ∎
Theorem 2.
Let $t_{1}\geq t_{2}$ be two non-negative integers. A code $\mathcal{C}$ can correct two bursts of $(t_{1},t_{2})$-DI if and only if it can correct one burst of $(t_{1},t_{2})$-DI together with one burst of $(t_{2},t_{1})$-DI.
Proof.
The proof follows a similar strategy to that of Theorem 1. An ECC $\mathcal{C}$ can correct two bursts of $(t_{1},t_{2})$-DI, that means, for any $\mathbf{x},\mathbf{y}\in\mathcal{C}$, we have
$\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})\cap\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{y})=\varnothing.$
Assume that $\mathcal{C}$ cannot correct one burst of $(t_{1},t_{2})$-DI together with one burst of $(t_{2},t_{1})$-DI. Then, there exist two different sequences $\mathbf{x},\mathbf{y}\in\mathcal{C}$ such that
$\mathcal{B}_{(t_{1},t_{2}),(t_{2},t_{1})}^{DI}(\mathbf{x})\cap\mathcal{B}_{(t_{1},t_{2}),(t_{2},t_{1})}^{DI}(\mathbf{y})\neq\varnothing,$
where $\mathcal{B}_{(t_{1},t_{2}),(t_{2},t_{1})}^{DI}(\mathbf{x})$ represents the set of sequences obtained from $\mathbf{x}$ by one burst of $(t_{1},t_{2})$-DI together with one burst of $(t_{2},t_{1})$-DI. That is to say, we can find a sequence $\mathbf{s}\in\Sigma_{2}^{n}$ such that
$\mathbf{s}\in\mathcal{B}_{(t_{1},t_{2}),(t_{2},t_{1})}^{DI}(\mathbf{x})\cap\mathcal{B}_{(t_{1},t_{2}),(t_{2},t_{1})}^{DI}(\mathbf{y}).$
Suppose that $\mathbf{s}$ is obtained from $\mathbf{x}$ by deleting $(x_{i_{1}},x_{i_{1}+1},\dots,x_{i_{1}+t_{1}-1})$ and $(x_{i_{2}},x_{i_{2}+1},\dots,x_{i_{2}+t_{2}-1})$, where $i_{2}\geq i_{1}+t_{1}$, and inserting $(b_{1},b_{2},\dots,b_{t_{2}})$ and $(b^{\prime}_{1},b^{\prime}_{2},\dots,b^{\prime}_{t_{1}})$ at the $i_{1}$-th and $i_{2}$-th components of $\mathbf{x}$, then we have
$\mathbf{s}=$ $x_{1},\dots,x_{i_{1}-1},b_{1},b_{2},\dots,b_{t_{2}},x_{i_{1}+t_{1}},\dots,x_{i_{2}-1},$
$b^{\prime}_{1},b^{\prime}_{2},\dots,b^{\prime}_{t_{1}},x_{i_{2}+t_{2}},\dots,x_{n}.$ (11)
Meanwhile, if $\mathbf{s}$ is obtained from $\mathbf{y}$ by deleting $(y_{j_{1}},y_{j_{1}+1},$ $\dots,y_{j_{1}+t_{1}-1})$ and $(y_{j_{2}},y_{j_{2}}+1,\dots,y_{j_{2}+t_{2}-1})$, where $j_{2}\geq j_{1}+t_{1}$, and inserting $(c_{1},c_{2},\dots,c_{t_{2}})$ and $(c^{\prime}_{1},c^{\prime}_{2},\dots,c^{\prime}_{t_{1}})$ at the $j_{1}$-th and $j_{2}$-th components of $\mathbf{y}$, then we have
$\mathbf{s}=$ $y_{1},\dots,y_{j_{1}-1},c_{1},c_{2},\dots,c_{t_{2}},y_{j_{1}+t_{1}},\dots,y_{j_{2}-1},$
$c^{\prime}_{1},c^{\prime}_{2},\dots,c^{\prime}_{t_{1}},y_{j_{2}+t_{2}},\dots,y_{n}.$ (12)
In the sequel, we only consider the case of $i_{1}+t_{1}\leq i_{2}\leq j_{1}-t_{2}\leq j_{2}-t_{1}-t_{2}$ to show the main idea of the proof. By comparing (III-A) and (III-A), we obtain
[…]
Therefore, by deleting $(x_{i_{1}},x_{i_{1}+1},\dots,x_{i_{1}+t_{1}-1})$ and $(x_{j_{2}+t_{2}-t_{1}},\dots,x_{j_{2}+t_{2}-1})$ from $\mathbf{x}$, and then inserting $(y_{i_{1}},\dots,y_{i_{1}+t_{2}-1})$ and $(y_{j_{2}},\dots,y_{j_{2}+t_{2}-1})$ at the $i_{1}$-th and $j_{2}+t_{2}-t_{1}$-th components of $\mathbf{x}$, respectively, we can obtain
$\mathbf{x}^{\prime\prime\prime}=(x_{1},\dots,x_{i_{1}-1},y_{i_{1}},\dots,y_{i_{1}+t_{2}-1},x_{i_{1}+t_{1}},\dots,$
$x_{j_{2}+t_{2}-t_{1}-1},y_{j_{2}},\dots,y_{j_{2}+t_{2}-1},x_{j_{2}+t_{2}},\dots,x_{n}).$
By deleting $(y_{i_{2}+t_{2}-t_{1}},\dots,y_{i_{2}+t_{2}-1})$ and $(y_{j_{1}},\dots,y_{j_{1}+t_{1}-1})$ from $\mathbf{y}$, and then inserting $(x_{i_{2}},\dots,x_{i_{2}+t_{2}-1})$ and $(x_{j_{1}},\dots,$ $x_{j_{1}+t_{2}-1})$ at the $i_{2}+t_{2}-t_{1}$-th and $j_{1}$-th components of $\mathbf{y}$, respectively, we obtain
$\mathbf{y}^{\prime\prime\prime}=($ $y_{1},\dots,y_{i_{2}+t_{2}-t_{1}-1},x_{i_{2}},\dots,x_{i_{2}+t_{2}-1},y_{i_{2}+t_{2}},\dots,$
$y_{j_{1}-1},x_{j_{1}},\dots,x_{j_{1}+t_{2}-1},y_{j_{1}+t_{1}},\dots,y_{n}).$
Following the equalities in (18), $\mathbf{x}^{\prime\prime\prime}=\mathbf{y}^{\prime\prime\prime}$ is clear. This implies that
$\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})\cap\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{y})\neq\varnothing,$
which contradicts with the fact that $\mathcal{C}$ is an ECC for correcting two bursts of $(t_{1},t_{2})$-DI.
The proof of the case where $\mathbf{s}$ is obtained from $\mathbf{y}$ by deleting $(y_{j_{1}},y_{j_{1}+1},\dots,y_{j_{1}+t_{2}-1})$ and $(y_{j_{2}},y_{j_{2}}+1,\dots,y_{j_{2}+t_{1}-1})$ with $j_{2}\geq j_{1}+t_{2}$, and inserting $(c_{1},c_{2},\dots,c_{t_{1}})$ and $(c^{\prime}_{1},c^{\prime}_{2},\dots,c^{\prime}_{t_{2}})$ at the $j_{1}$-th and $j_{2}$-th components of $\mathbf{y}$, respectively, is analogous to that of the previous case, thus completing the proof. ∎
The following corollary follows immediately from an argument similar to those of Theorems 1 and 2.
Corollary 1.
Let $t_{1}\geq t_{2}$ be two non-negative integers. A code $\mathcal{C}$ can correct $m$ bursts of $(t_{1},t_{2})$-DI if and only if it can correct $m_{1}$ bursts of $(t_{1},t_{2})$-DI together with $m_{2}$ bursts of $(t_{2},t_{1})$-DI as long as $m_{1}+m_{2}=m$.
III-B The Bounds on the Cardinality of Two Bursts of $(t_{1},t_{2})$-DI ECCs
In this subsection, we develop both lower and upper bounds on the size of ECCs that can correct two bursts of $(t_{1},t_{2})$-DI errors. The lower bound is an extension of two bursts of exactly $t_{1}$-deletion ECCs proposed in [35], corresponding to two bursts of $(t_{1},0)$-DI ECCs. On the other hand, the upper bound extends the result in [32], which considers one burst of $(t_{1},t_{2})$-DI ECCs.
We denote the cardinality of binary two bursts of $(t_{1},t_{2})$-DI ECC $\mathcal{C}$ with length $n$ satisfying $n>2t_{1}$ by $N_{\mathcal{C}}$.
Theorem 3 (Lower bound on the code size).
For any given $t_{1}$ and $t_{2}$, there exists a two bursts of $(t_{1},t_{2})$-DI ECC $\mathcal{C}\subseteq\Sigma_{2}^{n}$ with
$N_{\mathcal{C}}\geq\frac{2^{n-2t_{1}-2t_{2}}}{\binom{n-2t_{1}+2}{2}\binom{n-2t_{1}+2t_{2}+1}{2}}$.
Proof.
We define a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with vertex set $\mathcal{V}=\Sigma_{2}^{n}$, where two distinct vertices $\mathbf{v},\mathbf{v}^{\prime}\in\mathcal{V}$ are connected by an edge $(\mathbf{v},\mathbf{v}^{\prime})\in\mathcal{E}$ if and only if
$\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{v})\cap\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{v}^{\prime})\neq\varnothing.$
This adjacency relation is denoted compactly as $\mathbf{v}\sim\mathbf{v}^{\prime}$.
The independence number of $\mathcal{G}$, denoted by $N(\mathcal{V},\mathcal{E})$, represents the maximum cardinality of an independent set in $\mathcal{G}$. A code $\mathcal{C}\subseteq\Sigma_{2}^{n}$ is an ECC capable of correcting two bursts of $(t_{1},t_{2})$-DI if and only if $\mathcal{C}$ forms an independent set in $\mathcal{G}$. Consequently, the maximum code size satisfies $N_{\mathcal{C}}=N(\mathcal{V},\mathcal{E})$.
For a given vertex $\mathbf{x}\in\mathcal{V}$, let $\deg(\mathbf{x})$ denote its degree in $\mathcal{G}$, i.e., the number of vertices $\mathbf{y}\in\mathcal{V}$ such that $\mathbf{x}\sim\mathbf{y}$. Applying the probabilistic method from [1], we derive the following lower bound on the code size,
$N_{\mathcal{C}}\geq\sum_{\mathbf{x}\in\Sigma_{2}^{n}}\frac{1}{\deg(\mathbf{x})+1}.$
To enumerate the compatible pairs $(\mathbf{x},\mathbf{y})$, we observe that any $\mathbf{y}$ satisfying $\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})\cap\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{y})\neq\varnothing$ can be obtained via the following combinatorial operations: (1) Delete two non-overlapping substrings of length $t_{1}$ from $\mathbf{x}$, yielding an intermediate string $\mathbf{x}^{\prime}\in\Sigma_{2}^{n-2t_{1}}$, and subsequently insert two substrings of length $t_{2}$ at the deleted positions, resulting in $\mathbf{x}^{\prime\prime}\in\Sigma_{2}^{n-2t_{1}+2t_{2}}$. (2) Insert two substrings of length $t_{1}$ into $\mathbf{x}^{\prime\prime}$ and then delete two substrings of length $t_{2}$ immediately preceding the inserted positions to obtain $\mathbf{y}$. This construction yields the upper bound
$\deg(\mathbf{x})+1\leq\binom{n-2t_{1}+2}{2}\binom{n-2t_{1}+2t_{2}+1}{2}2^{2(t_{1}+t_{2})},$
thus completing the proof. ∎
Theorem 4 (Upper bound on the code size).
For any given two non-negative integers $t_{1}\geq t_{2}\geq 1$, the cardinality of an ECC capable of correcting two bursts of $(t_{1},t_{2})$-DI satisfies
$N_{\mathcal{C}}\leq{2^{n-2t_{1}+2t_{2}}\over{\max(a_{1},a_{2})}},$
where
$a_{1}=2^{2t_{2}-2}\left({n-2t_{1}+2\choose 2}+1\right),$
$a_{2}=2^{2t_{1}-2}\left({n-2t_{2}+2\choose 2}+1\right).$
Proof.
For any $\mathbf{x}\in\mathcal{C}$, assume that $\mathbf{x}^{\prime}$ and $\mathbf{x}^{\prime\prime}$ are obtained from $\mathbf{x}$ by two bursts of $(t_{1},t_{2})$-DI at the $i_{1}$-th, $i_{2}$-th and $i^{\prime}_{1}$-th, $i^{\prime}_{2}$-th components of $\mathbf{x}$, respectively, where $1\leq i_{1}\leq i_{2}-t_{1}\leq n-2t_{1}+1,1\leq i^{\prime}_{1}\leq i^{\prime}_{2}-t_{1}\leq n-2t_{1}+1$ and $\{i_{1},i_{2}\}\neq\{i^{\prime}_{1},i^{\prime}_{2}\}$. Following the definition of one burst of $(t_{1},t_{2})$-DI, it is clear that
$x^{\prime}_{i_{1}}\neq x_{i_{1}},x^{\prime}_{i_{2}+t_{2}-t_{1}}\neq x_{i_{2}},$
and
$x^{\prime\prime}_{i^{\prime}_{1}}\neq x_{i^{\prime}_{1}},x^{\prime\prime}_{i^{\prime}_{2}+t_{2}-t_{1}}\neq x_{i^{\prime}_{2}}.$
Without loss of the generality, let $i_{1}\leq i^{\prime}_{1}$ and $i_{2}\leq i^{\prime}_{2}$. Then, we can obtain $\mathbf{x}^{\prime}\neq\mathbf{x}^{\prime\prime}$ since $x^{\prime\prime}_{i_{1}}=x_{i_{1}}$ when $i_{1}<i^{\prime}_{1}$ or since $x^{\prime\prime}_{i_{2}+t_{2}-t_{1}}=x_{i_{2}}$ when $i_{1}=i^{\prime}_{1}$ and $i_{2}<i^{\prime}_{2}$. Consequently, for different index pairs $\{i_{1},i_{2}\}\neq\{i^{\prime}_{1},i^{\prime}_{2}\}$, the resulting sequences are not identical.
Define
$\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x},i_{1},i_{2})\triangleq$
$\left\{\begin{array}[]{lll}\mathbf{x}^{\prime}\in\Sigma_{2}^{n-2t_{1}+2t_{2}}:\mathbf{x}^{\prime}_{[n-2t_{1}+2t_{2}]\backslash(\mathcal{I}^{\prime}_{1}\cup\mathcal{I}^{\prime}_{2})}=\mathbf{x}_{[n]\backslash(\mathcal{I}_{1}\cup\mathcal{I}_{2})},\\ \hskip 73.97733pt\text{~and~}x^{\prime}_{i_{1}}\neq x_{i_{1}},x^{\prime}_{i_{2}-t_{1}+t_{2}}\neq x_{i_{2}},\\ \hskip 68.28644pt\text{~if~}1\leq i_{1}\leq i_{2}-t_{1}\leq n-2t_{1}+1,\\ \mathbf{x}^{\prime}\in\Sigma_{2}^{n-2t_{1}+2t_{2}}:\mathbf{x}^{\prime}_{[n-2t_{1}+1]}=\mathbf{x}_{[n-2t_{1}+1]},\\ \hskip 68.28644pt\text{~if~}i_{1}=n-2t_{1}+2,i_{2}=n-t_{1}+2.\end{array}\right$.
where $\mathcal{I}^{\prime}_{1}=[i_{1},i_{1}+t_{2}-1],\mathcal{I}^{\prime}_{2}=[i_{2}-t_{1}+t_{2},i_{2}-t_{1}+2t_{2}-1],\mathcal{I}_{1}=[i_{1},i_{1}+t_{1}-1]$ and $\mathcal{I}_{2}=[i_{2},i_{2}+t_{1}-1]$.
For any $1\leq i_{1}\leq i_{2}-t_{1}\leq n-2t_{1}+2,1\leq i^{\prime}_{1}\leq i^{\prime}_{2}-t_{1}\leq n-2t_{1}+2$ with $\{i_{1},i_{2}\}\neq\{i^{\prime}_{1},i^{\prime}_{2}\}$, it is obvious that
$\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x},i_{1},i_{2})\subseteq\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})$
and
$\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x},i_{1},i_{2})\cap\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x},i^{\prime}_{1},i^{\prime}_{2})=\varnothing.$
Then if
$\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})\subseteq\bigcup_{i_{1}=1}^{n-2t_{1}+2}\bigcup_{i_{2}=i_{1}+t_{1}}^{n-t_{1}+2}\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x},i_{1},i_{2}),$ (19)
we can state that $\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x},i_{1},i_{2})$ with $i_{1}\in[n-2t_{1}+2],i_{2}\in[i_{1}+t_{1},n-t_{1}+2]$ form a partition of $\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})$.
In what follows, we prove that the equality in (19) holds. Actually, for any $\mathbf{x}^{\prime}\in\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})$, if there exist two indices $i_{1}\in[n-2t_{1}+1]$ and $i_{2}\in[i_{1}+t_{1},n-t_{1}+1]$ such that $\mathbf{x}^{\prime}_{[n-2t_{1}+1]}\neq\mathbf{x}_{[n-2t_{1}+1]}$, then $\mathbf{x}^{\prime}\in\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x},i_{1},i_{2})$, otherwise, $\mathbf{x}^{\prime}\in\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x},n-2t_{1}+2,n-t_{1}+2)$.
Therefore,
$\begin{aligned} \left|\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})\right|=&\sum_{i_{1}=1}^{n-2t_{1}+1}\sum_{i_{2}=i_{1}+t_{1}}^{n-t_{1}+1}\left|\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x},i_{1},i_{2})\right|\\ &+\left|\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x},n-2t_{1}+2,n-t_{1}+2)\right|\\ =&2^{2t_{2}-2}\sum_{i_{1}=1}^{n-2t_{1}+1}(n-2t_{1}-i_{1}+2)+2^{2t_{2}-2}\\ =&2^{2t_{2}-2}\left({n-2t_{1}+2\choose 2}+1\right).\end{aligned}$
Since $\mathcal{C}$ is a two bursts of $(t_{1},t_{2})$-DI ECC, we have
$\left|\bigcup_{\mathbf{x}\in\mathcal{C}}\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})\right|=\sum_{\mathbf{x}\in\mathcal{C}}\left|\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})\right|$
$=$ $\left|\mathcal{C}\right|\cdot 2^{2t_{2}-2}\left({n-2t_{1}+2\choose 2}+1\right)$
$\leq$ $2^{n-2t_{1}+2t_{2}}.$
Moreover, based on Theorem 1, $\mathcal{C}$ is also a two bursts of $(t_{2},t_{1})$-DI ECC, we have
$\left|\bigcup_{\mathbf{x}\in\mathcal{C}}\mathcal{B}_{2,(t_{2},t_{1})}^{DI}(\mathbf{x})\right|=\sum_{\mathbf{x}\in\mathcal{C}}\left|\mathcal{B}_{2,(t_{2},t_{1})}^{DI}(\mathbf{x})\right|$
$=$ $\left|\mathcal{C}\right|\cdot 2^{2t_{1}-2}\left({n-2t_{2}+2\choose 2}+1\right)$
$\leq$ $2^{n-2t_{1}+2t_{2}}.$
Denoting $a_{1}=2^{2t_{2}-2}\left({n-2t_{1}+2\choose 2}+1\right)$ and $a_{2}=2^{2t_{1}-2}\left({n-2t_{2}+2\choose 2}+1\right)$, the equations above imply that
$|\mathcal{C}|\leq\frac{2^{n-2t_{1}+2t_{2}}}{\max(a_{1},a_{2})},$
completing the proof. ∎
Based on Theorems 3 and 4, the redundancy of two bursts of $(t_{1},t_{2})$-DI ECCs satisfies
$\log(\max(a_{1},a_{2}))+2t_{1}-2t_{2}\leq~r(\mathcal{C})$
$\leq$ $\log\left(\binom{n-2t_{1}+2}{2}\binom{n-2t_{1}+2t_{2}+1}{2}\right)+2t_{1}+2t_{2}.$
Furthermore, we have the following corollary.
Corollary 2.
For any ECC $\mathcal{C}\in\Sigma_{2}^{n}$ capable of correcting two bursts of $(t_{1},t_{2})$-DI, the redundancy of $\mathcal{C}$ satisfies $2\log n+O(1)\leq r(\mathcal{C})\leq 4\log n+o(\log n)$.
Remark 1.
When $t_{1}=t_{2}=1$, two bursts of $(1,1)$-DI coincide with exactly two substitution errors. In this case, the Hamming ball of size ${n\choose 2}+1$ is resulted, which is consistent with the Equation (20).
Remark 2.
The upper and lower bounds on the code size of two bursts of $(t_{1},t_{2})$-DI ECCs can be naturally extended to the case of $m$ bursts with $m\geq 2$, yielding the redundancy bounds of a code $\mathcal{C}\in\Sigma_{2}^{n}$ capable of correcting $m$ bursts of $(t_{1},t_{2})$-DI errors $m\log n+O(1)\leq r(\mathcal{C})\leq 2m\log n+o(\log n)$.
IV Two bursts of $(t_{1},t_{2})$-DI ECCs by Syndrome Compression Techniques
In view of the equivalence results established in Section III, in what follows, we restrict our attention to two bursts of $(t_{1},t_{2})$-DI errors with $t_{1}\geq t_{2}$. Under this setting, we develop in this section constructions of two bursts of $(t_{1},t_{2})$-DI ECCs by using the syndrome compression technique proposed in [24]. Although the applicability of syndrome compression technique to the construction of such codes is relatively straightforward, we provide the detailed construction here to lay the groundwork for the next section and to facilitate a comparison in terms of computational complexity.
Before presenting the construction, we first review a function capable of correcting multiple deletions, insertions and substitutions together.
Lemma 2.
(Section V, [23]) For any binary sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})\in\Sigma_{2}^{n}$, there exists a function $h:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{[(t^{2}+1)(2t^{2}+1)+2t^{2}(t-1)]\log n+o(\log n)}$, such that given $h(\mathbf{x})$ and $\mathbf{x}^{\prime}$ obtained from $\mathbf{x}$ by $t_{1}$-deletions, $t_{2}$-insertions and $t_{3}$-substitutions satisfying $t_{1}+t_{2}+t_{3}=t$, one can uniquely recover $\mathbf{x}$.
For any $\mathbf{x}\in\Sigma_{2}^{n}$, define
$\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})\triangleq\Big\{\mathbf{x}^{\prime}\in\Sigma_{2}^{n}:\mathbf{x}^{\prime}\neq\mathbf{x},\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x}^{\prime})\cap\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})\neq\varnothing\Big\},$
where $t_{1}$ and $t_{2}$ are two constants with respect to $n$.
In fact, each codeword in $\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})$ can be obtained via the following steps.
  • Delete two substrings of length $t_{1}$ from $\mathbf{x}$, and then insert two substrings of length $t_{2}$ at the corresponding deletion positions to obtain a sequence $\mathbf{y}\in\Sigma_{2}^{n-2t_{1}+2t_{2}}$, where the first and last symbols of each inserted substring are different from the first and last symbols of the deleted substring at the same position, respectively. There are at most ${n-2t_{1}+2\choose 2}2^{2t_{2}-4}$ and $n-2t_{1}+2\choose 2$} possibilities for $\mathbf{y}$ when $t_{2}\geq 2$ and $t_{2}=1$, respectively.
  • Insert two substring of length $t_{1}$ and then delete two substrings of length $t_{2}$ from $\mathbf{y}$ following the insertion positions to obtain a sequence $\mathbf{x}^{\prime}\in\Sigma_{2}^{n}$, where the first and last symbols of each inserted substring are different from the first and last symbols of the deleted substring at the same position, respectively. There are at most ${n-2t_{1}+2t_{2}+1\choose 2}\cdot 2^{2t_{1}-4}$ and $n-2t_{1}+2t_{2}+1\choose 2$} possibilities for $\mathbf{x}^{\prime}$ when $t_{1}\geq 2$ and $t_{1}=1$, respectively.
This results in the cardinality of $\mathcal{N}^{DI}_{2,(t_{1},t_{2})}$ being
$\begin{aligned} \left|\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})\right|=&{n-2t_{1}+2\choose 2}{n-2t_{1}+2t_{2}+1\choose 2}\cdot 2^{2(t_{1}+t_{2})-8}\\ \leq&n^{4}2^{2(t_{1}+t_{2})}\end{aligned}$
when $t_{1},t_{2}\geq 2$, and
$\begin{aligned} \left|\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})\right|=&{n-2t_{1}+2\choose 2}{n-2t_{1}+2t_{2}+1\choose 2}\cdot 2^{2t_{1}-4}\\ \leq&n^{4}2^{2t_{1}-4}\end{aligned}$
when $t_{1}>t_{2}=1$.
Construction 1.
For any $\mathbf{x}\in\Sigma_{2}^{n}$, let $h(\mathbf{x})$ be the function proposed in Lemma 2 for correcting $2t_{1}$-deletions and $2t_{2}$-insertions, and $\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})$ be the set defined in (21). Denote
$f_{1}(\mathbf{x})=\Big(a(\mathbf{x}),h(\mathbf{x})\bmod a(\mathbf{x})\Big),$ (22)
where $a(\mathbf{x})$ is a non-negative integer satisfying
$a(\mathbf{x})\leq 2^{\log{\left|\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})\right|+o(\log n)}}$
and $a(\mathbf{x})\notin\{j:j\mid(h(\mathbf{x})-h(\mathbf{x}^{\prime}))\}$ for any $\mathbf{x}^{\prime}\in\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})$.
For a fixed parameter $f_{1}\in[0,2|a(\mathbf{x})|)$, define a code by
$\mathcal{C}_{1}\triangleq\Big\{\mathbf{x}\in\Sigma_{2}^{n}:f_{1}(\mathbf{x})=f_{1}\Big\}.$
We can state the following conclusion.
Theorem 5.
The code $\mathcal{C}_{1}$ generated in Construction 1 can correct two bursts of $(t_{1},t_{2})$-DI in a binary sequence of length $n$ with at most $8\log n+o(\log n)$ bits of redundancy.
Proof.
Since the function $h(\mathbf{x})$ proposed in (22) can correct $2t_{1}$-deletions and $2t_{2}$-insertions, it can correct two bursts of $(t_{1},t_{2})$-DI. Viewing $h(\mathbf{x})$ as the binary representation of a nonnegative integer, by brute force search, one can find a positive integer $a(\mathbf{x})$ in time
$2^{\log{\left|\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})\right|+o(\log n)}}\leq O(n^{4}).$
Since $h(\mathbf{x})\neq h(\mathbf{x}^{\prime\prime})$, then $f_{1}(\mathbf{x})\neq f_{1}(\mathbf{x}^{\prime\prime})$ for $\mathbf{x}^{\prime\prime}\in\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})$. Equivalently, the original sequence $\mathbf{x}$ can be uniquely recovered by $f_{1}(\mathbf{x})$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})$.
The redundancy of $f_{1}(\mathbf{x})$ in (22) is
$2\left|a(\mathbf{x})\right|=$ $2\log\left|\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})\right|+o(\log n)$
$\leq$ $8\log n+o(\log n),$
completing the proof. ∎
Remark 3.
The code $\mathcal{C}_{1}$ generated in Construction 1 can correct one burst of $(t_{1},t_{2})$-DI with at most $4\log n+o(\log n)$ bits of redundancy.
While the functions derived from the syndrome compression technique are indeed capable of correcting errors in a sequence, applying this technique directly to binary sequences typically results in prohibitively high computational complexity, making their practical implementation challenging. To address this issue, we construct ECCs capable of correcting two bursts of $(t_{1},t_{2})$-DI by leveraging an matrix-based representation of the sequences in the next section.
As a basis for the next section, we note that the following conclusion can be obtained by an argument similar to that used in Construction 1.
For any $\mathbf{x}\in\Sigma_{2}^{n}$, define
$\mathcal{N}^{DS}_{2,(1,t)}(\mathbf{x})\triangleq\Big\{\mathbf{x}^{\prime}\in\Sigma_{2}^{n}:\mathbf{x}^{\prime}\neq\mathbf{x},\mathcal{B}_{2,(1,t)}^{DS}(\mathbf{x}^{\prime})\cap\mathcal{B}_{2,(1,t)}^{DS}(\mathbf{x})\neq\varnothing\Big\},$
where $t\geq 1$ is a constant with respect to $n$. The cardinality of $\mathcal{N}^{DS}_{2,(1,t)}$ is
$\begin{aligned} \left|\mathcal{N}^{DS}_{2,(1,t)}(\mathbf{x})\right|=&{n-2t\choose 2}{n-2t+1\choose 2}\Big(\sum_{i=0}^{t-1}{t-1\choose i}\Big)^{4}\\ \leq&n^{4}2^{4t-4}.\end{aligned}$
Therefore, the following Lemma holds.
Lemma 3.
For any binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$, there exists a function $f:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{n+8\log n+o(\log n)}$, such that given $\mathbf{x}^{\prime}\in\mathcal{B}_{2,(1,t)}^{DS}(\mathbf{x})$ and $f(\mathbf{x})$, one can uniquely recover $\mathbf{x}$.
V Two Bursts of $(t_{1},t_{2})$-DI ECCs
In this section, we present constructions of binary ECCs capable of correcting two bursts of $(t_{1},t_{2})$-DI, where $t_{1}\geq t_{2}$ are two positive integers.
V-A Two Bursts of $(t,t)$-DI ECCs
In this subsection, a construction is proposed to correct two bursts of $(t,t)$-DI for binary sequences, i.e., $t_{1}=t_{2}=t$. In fact, two bursts of $(t,t)$-DI in a binary sequence is the same as two bursts of $t$-substitutions with the first and the last symbols are substituted necessarily, which can be converted to two bursts of $(2,2)$-DI of $2^{t-1}$-ary sequences following the results in [32].
For any binary sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})\in\Sigma_{2}^{n}$, assume that $(t-1)\mid n$, otherwise we append zeros at the end of each sequence such that its length is the smallest integer greater than $n$ and divisible by t1t-1. Then, when $t\geq 2$, the binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$ can be represented by a $(t-1)\times(n/(t-1))$ binary matrix as follows,
$X_{t-1}=\left(\begin{array}[]{cccccc}x_{1}&x_{t}&\cdots&x_{n-t+2}\\ x_{2}&x_{t+1}&\cdots&x_{n-t+3}\\ \vdots&\vdots&\ddots&\vdots\\ x_{t-1}&x_{2t-2}&\cdots&x_{n}\end{array}\right).$
Denote $\tilde{\mathbf{x}}$ is a $2^{t-1}$-ary vector corresponding to $X_{t-1}$. Specifically, let
$\tilde{\mathbf{x}}=(\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{n\over{t-1}})\in\Sigma_{2^{t-1}}^{n\over{t-1}},$ (24)
where
$\tilde{x}_{i}=\sum_{j=1}^{t-1}2^{t-1-j}\cdot x_{(i-1)(t-1)+j},\quad i\in\left[{n\over{t-1}}\right].$
Lemma 4 (Theorem 3, [32]).
Let $\mathbf{x}\in\Sigma_{2}^{n}$ be a binary sequence of length $n$ and $\tilde{\mathbf{x}}$ be the corresponding $2^{t-1}$-ary vector defined in (24), where $t\geq 2$ is a positive integer. If a $(t,t)$-DI occurs in $\mathbf{x}$, then there is a $(2,2)$-DI in $\tilde{\mathbf{x}}$.
Theorem 6.
For any $\mathbf{x}\in\Sigma_{2}^{n}$ and an integer $t\geq 2$, two bursts of $(t,t)$-DI in $\mathbf{x}$ can be corrected by an $n/(t-1),n/(t-1)-8,d_{\mathrm{min}}=9]_{q}$ Generalized Reed-Solomon (GRS) code with at most $8\log n+O(1)$ bits of redundancy, where $q\geq n$.
Proof.
According to Lemma 4, it is clear that two bursts of $(t,t)$-DI in the original sequence $\mathbf{x}$ induce two bursts of $(2,2)$-DI in the transformed sequence $\tilde{\mathbf{x}}$. These errors manifest as two bursts of two symbol substitutions in $\tilde{\mathbf{x}}$, which together amount to a total of four symbol substitutions. Such error patterns can be corrected by using a GRS code with the minimum Hamming distance $9$.
Since $\tilde{\mathbf{x}}\in\Sigma_{2^{t-1}}^{n/(t-1)}$, the GRS code must be defined over a finite field $\mathbb{F}_{q}$ with $q\geq\max\big(2^{t-1},n/(t-1)\big)$. As $t\geq 2$ being a constant independent of $n$, it suffices to choose $q$ as the smallest prime satisfying $q\geq n$. Hence, an $n/(t-1),n/(t-1)-8,d_{\mathrm{min}}=9]_{q}$ GRS code with $q\geq n$ can be directly applied. As illustrated in [18], a GRS code over an alphabet of size $q\geq n$ for correcting $t$-symbol errors requires $2t\log q$ bits of redundancy. Consequently, correcting four substitution errors incurs a redundancy with at most $8\log n+O(1)$ bits. ∎
Remark 4.
When $t=1$, two bursts of $(1,1)$-DI in $\mathbf{x}\in\Sigma_{2}^{n}$ cause two substitutions in $\mathbf{x}$, then applying an $n,n-4,d_{\mathrm{min}}=5]_{q}$ GRS code with at most $4\log n+O(1)$ bits of redundancy is sufficient, where $q\geq n$.
V-B Two Bursts of $(t_{1},t_{2})$-DI ECCs
In this subsection, we consider the case of $t_{1}>t_{2}$ and address design of ECCs correcting two bursts of $(t_{1},t_{2})$-DI errors. By representing a sequence in an matrix formula, we can demonstrate that two bursts of $(t_{1},t_{2})$-DI cause two bursts of $(1,t^{\prime}-1)$-DS in each row of the matrix, where $t^{\prime}=\lceil t_{1}/(t_{1}-t_{2})\rceil$. Then, by locating the positions of the two bursts in the first row, the error positions of the original sequence can be approximately identified. Finally, two bursts of $(t_{1},t_{2})$-DI in the original sequence can be corrected following the function proposed in Construction 1 with reduced computational complexity.
For any binary sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})\in\Sigma_{2}^{n}$, assume that $(t_{1}-t_{2})\mid n$, otherwise we append zeros at the end of each sequence such that its length is the smallest integer greater than $n$ and divisible by $t_{1}-t_{2}$. Then, the binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$ can be represented by a $(t_{1}-t_{2})\times(n/(t_{1}-t_{2}))$ binary matrix as follows,
$\begin{aligned} X_{t_{1}-t_{2}}&=\left(\begin{array}[]{cccccc}X[1]\\ X[2]\\ \vdots\\ X[t_{1}-t_{2}]\end{array}\right)=\left(\begin{array}[]{cccccc}x_{1}&\cdots&x_{n-t_{1}+t_{2}+1}\\ x_{2}&\cdots&x_{n-t_{1}+t_{2}+2}\\ \vdots&\ddots&\vdots\\ x_{t_{1}-t_{2}}&\cdots&x_{n}\end{array}\right),\end{aligned}$
where $X[i],i\in[t_{1}-t_{2}],$ refers to the $i$-th row of $X_{t_{1}-t_{2}}$.
Denote $t^{\prime}=\lceil t_{1}/(t_{1}-t_{2})\rceil$ and
$\mathcal{M}=\left\{m\in[n]:m=a(t_{1}-t_{2})+1,a\in\left[0,{n\over t_{1}-t_{2}}\right)\right\}.$
The following fact can be obtained by analyzing the matrix representation of $\mathbf{x}$ in (25).
Observation 2.
For any binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$ and $i\in[n-t_{1}+1]$, suppose that one burst of $(t_{1},t_{2})$-DI occurs at the $i$-th component of $\mathbf{x}$. Then, one burst of $(1,t^{\prime}-1)$-DS occurs at the $b_{1}$-th component of $X[1]$ when $i\in\mathcal{M}$, where $b_{1}=\lceil i/(t_{1}-t_{2})\rceil$, or the $b_{2}$-th component of $X[1]$ when i[$\i\in[n]\backslash\mathcal{M}$, where $b_{2}\in[\lceil i/(t_{1}-t_{2})\rceil+1,\lceil i/(t_{1}-t_{2})\rceil+t^{\prime})$. For other rows of $X_{t_{1}-t_{2}}$, the starting position of the errors is at most $t^{\prime}$ columns before the starting position of the errors in the first row.
Example 1.
For given $n=18,t_{1}=10,t_{2}=7$, we have $t^{\prime}=4$. Assume that
$\mathbf{x}=(0,1,1,0,0,0,0,0,1,1,0,1,1,0,1,0,1,1),$
if $(0,0,0,0,0,1,1,0,1,1)$ is deleted and $(1,0,1,0,1,1,0)$ is inserted at the fourth component of $\mathbf{x}$, that is,
$X_{3}=\left(\begin{array}[]{ccccccccccc}0&\underline{0}&\underline{0}&\underline{1}&\underline{1}&0\\ 1&\underline{0}&\underline{0}&\underline{0}&0&1\\ 1&\underline{0}&\underline{1}&\underline{1}&1&1\end{array}\right)\rightarrow\left(\begin{array}[]{ccccccccccc}0&\bar{1}&\bar{0}&\bar{0}&0\\ 1&\bar{0}&\bar{1}&0&1\\ 1&\bar{1}&\bar{1}&1&1\end{array}\right),$
where $\underline{i}$ is the deleted bit and $\bar{i}$ is the inserted bit in $\mathbf{x}$, $i\in\{0,1\}$, then there is one burst of $(1,3)$-DS at the second component of $X[1]$ and the errors in $X[2]$ and $X[3]$ start from the third and second components, respectively.
If $(0,0,0,0,1,1,0,1,1,0)$ is deleted and the following cases occur,
$X_{3}=\begin{pmatrix}0&0&\underline{0}&\underline{1}&\underline{1}&0\\ 1&\underline{0}&\underline{0}&\underline{0}&\underline{0}&1\\ 1&\underline{0}&\underline{1}&\underline{1}&1&1\end{pmatrix}$
[…]
then one burst of $(1,2)$-DS, one burst of $(1,1)$-DS and $1$-deletion occur at the third, fourth, and fifth components of $X[1]$ in (26), (27) and (28), respectively, and the errors in $X[2]$ and $X[3]$ start at most three columns before the starting position of the error in $X[1]$.
In this section, we focus on binary sequences whose representation under $X_{t_{1}-t_{2}}$ has a $d$-regular first row. The following lemma presents an explicit encoding function that transforms arbitrary binary sequences into ones whose first row of the matrix representation is $d$-regular.
Lemma 5.
For any binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$, there exists a one-to-one mapping $\mathcal{E}:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{n+1}$, such that for $\mathbf{x}^{\prime}=\mathcal{E}(\mathbf{x})$, if $X^{\prime}_{t_{1}-t_{2}}$ is the matrix representation of $\mathbf{x}^{\prime}$, then $X^{\prime}[1]$ is a $d$-regular sequence.
Proof.
Following (25), we apply the encoding function $\mathrm{RegEnc}$ in Lemma 1 to $X[1]$ such that
$\mathrm{RegEnc}(X[1])=(u_{1},u_{2},\cdots,u_{n/(t_{1}-t_{2})},u_{n/(t_{1}-t_{2})+1})$
is a $d$-regular sequence. Then, let
$\begin{aligned} X^{\prime}_{t_{1}-t_{2}}=\left(\begin{array}[]{cccccc}u_{1}&\cdots&u_{n/(t_{1}-t_{2})}&u_{n/(t_{1}-t_{2})+1}\\ x_{2}&\cdots&x_{n-t_{1}+t_{2}+2}&0\\ \vdots&\ddots&\vdots&\vdots\\ x_{t_{1}-t_{2}}&\cdots&x_{n}&0\end{array}\right),\end{aligned}$
[…]
Letting $\mathbf{x}^{\prime}=\mathcal{E}(\mathbf{x})$, the existence of such a function is guaranteed. Since $\mathrm{RegEnc}$ is a one-to-one mapping and all remaining symbols are copied unchanged, the mapping $\mathcal{E}$ is injective. Moreover, the function $\mathcal{E}$ increases the sequence length by only one bit. Finally, as $\mathrm{RegEnc}$ can be computed in near-linear time, the same holds for $\mathcal{E}$. ∎
Remark 5.
In Lemma 1, the encoding function $\mathrm{RegEnc}$ can be computed with only one bit of redundancy by relying on a lookup table of polynomial size. This claims that the computational complexity can be reduced at the expense of increased storage complexity. To mitigate the storage overhead, one may alternatively employ $(\mathbf{p}=0011,\delta=64\lceil\log n\rceil)$-dense sequences in place of $d$-regular sequence in this paper, since the encoding of the former in [27] can be performed in $O(n\log n)$ time and does not require the use of a look-up table.
In the following, we consider two bursts of $(t_{1},t_{2})$-DI in a binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$. Recall that the two bursts under consideration in this subsection are mutually non-overlapping.
To proceed, we need the following lemma.
Lemma 6.
For a $d$-regular sequence $\mathbf{x}\in\Sigma_{2}^{n}$, if $\mathbf{x}^{\prime}$ is obtained from $\mathbf{x}$ by two bursts of (1,t1)(1,t-1)-DS, then one of the following conclusions holds.
  • a)
    There are two distinct intervals $\mathcal{I}_{1},\mathcal{I}_{2}\subseteq[n]$ with $|\mathcal{I}_{1}|,|\mathcal{I}_{2}|\leq d\log n$, such that one burst of (1,t1)(1,t-1)-DS occurs in $\mathbf{x}_{\mathcal{I}_{1}}$ and the other in $\mathbf{x}_{\mathcal{I}_{2}}$, respectively.
  • b)
    There is an interval $\mathcal{I}\subseteq[n]$ with $|\mathcal{I}|\leq 3d\log n+2t$, such that two bursts of (1,t1)(1,t-1)-DS occur in $\mathbf{x}_{\mathcal{I}}$.
Proof.
Assume that $\mathbf{x}^{\prime}$ is obtained from $\mathbf{x}$ by two bursts of (1,t1)(1,t-1)-DS at the $i_{1}$-th and $i_{2}$-th components, where $1\leq i_{1}\leq i_{2}-t\leq n-2t+1$. If there exist two indices $j_{1},j_{2}$ such that $\mathbf{x}^{\prime}$ can also be obtained from $\mathbf{x}$ by two bursts of (1,t1)(1,t-1)-DS at the $j_{1}$-th and $j_{2}$-th components, where $1\leq j_{1}\leq j_{2}-t\leq n-2t+1$. In the sequel, we analyze scenarios in which the four burst positions are either disjoint or overlapping. Firstly, we consider the former by the following three cases, due to symmetry.
Case i) $i_{1}+t\leq j_{1}\leq i_{2}-t\leq j_{2}-2t$. In this case, we have
$\mathbf{x}^{\prime}=$ $x_{1}\cdots x_{i_{1}-1}\overline{x_{i_{1}+1}\cdots x_{i_{1}+t-1}}~x_{i_{1}+t}~~\cdots~~x_{j_{1}}~x_{j_{1}+1}\cdots$
$x_{j_{1}+t-1}x_{j_{1}+t}\cdots x_{i_{2}-1}\overline{x_{i_{2}+1}\cdots x_{i_{2}+t-1}}~x_{i_{2}+t}\cdots~x_{j_{2}}$~~
$x_{j_{2}+1}\cdots x_{j_{2}+t-1}x_{j_{2}+t}\cdots x_{n}$
$=$ $x_{1}\cdots x_{i_{1}-1}~~x_{i_{1}}~\cdots x_{i_{1}+t-2}x_{i_{1}+t-1}\cdots x_{j_{1}-1}\overline{x_{j_{1}+1}\cdots}$
$\overline{x_{j_{1}+t-1}}x_{j_{1}+t}\cdots x_{i_{2}-1}~~x_{i_{2}}~\cdots x_{i_{2}+t-2}x_{i_{2}+t-1}\cdots x_{j_{2}-1}$
$\overline{x_{j_{2}+1}\cdots x_{j_{2}+t-1}}x_{j_{2}+t}\cdots x_{n},$
where $\overline{x_{i+1}\cdots x_{i+t-1}}$ means that at most t1t-1 substitutions occur in $(x_{i+1}\cdots x_{i+t-1}),i\in\{i_{1},i_{2},j_{1},j_{2}\}$.
Viewing $\mathbf{x}^{\prime}$, it is impossible that $\overline{x_{i_{1}+1}\cdots x_{i_{1}+t-1}}=x_{i_{1}}\cdots x_{i_{1}+t-2}$ and $\overline{x_{i_{2}+1}\cdots x_{i_{2}+t-1}}=x_{i_{2}}\cdots x_{i_{2}+t-2}$ since $\overline{x}_{i_{1}+1}\neq x_{i_{1}}$ and $\overline{x}_{i_{2}+1}\neq x_{i_{2}}$ according to Definition 2. This indicates that, two bursts of (1,t1)(1,t-1)-DS in $\mathbf{x}$ degrades into two deletions in this case, i.e.,
$\mathbf{x}^{\prime}=$ $x_{1}\cdots x_{i_{1}-1}x_{i_{1}+1}\cdots~~x_{j_{1}}~x_{j_{1}+1}\cdots x_{i_{2}-1}x_{i_{2}+1}\cdots$~
$~x_{j_{2}}~~x_{j_{2}+1}\cdots x_{n}$
$=$ $x_{1}\cdots x_{i_{1}-1}~~x_{i_{1}}~\cdots x_{j_{1}-1}x_{j_{1}+1}\cdots x_{i_{2}-1}~~x_{i_{2}}$$\cdots$
$x_{j_{2}-1}x_{j_{2}+1}\cdots x_{n}.$
By comparing the symbols of the corresponding positions, we obtain
$x_{i_{1}}=x_{i_{1}+1}=\cdots=x_{j_{1}-1}=x_{j_{1}},$ (29)
$x_{i_{2}}=x_{i_{2}+1}=\cdots=x_{j_{2}-1}=x_{j_{2}},$
Following (29), $\mathbf{x}_{[i_{1},j_{1}]}$ and $\mathbf{x}_{[i_{2},j_{2}]}$ are two runs of $\mathbf{x}$. Denoting $\mathcal{I}_{1}=[i_{1},j_{1}]$ and $\mathcal{I}_{2}=[i_{2},j_{2}]$, we have $|\mathcal{I}_{1}|,|\mathcal{I}_{2}|\leq d\log n$ with Observation 1, then conclusion a) holds.
Case ii) $i_{1}+t\leq j_{1}\leq j_{2}-t\leq i_{2}-2t$. In this case, we have
$\mathbf{x}^{\prime}=$ $x_{1}\cdots x_{i_{1}-1}\overline{x_{i_{1}+1}\cdots x_{i_{1}+t-1}}~x_{i_{1}+t}~~\cdots~~x_{j_{1}}~x_{j_{1}+1}\cdots$
$x_{j_{1}+t-1}x_{j_{1}+t}\cdots x_{j_{2}-1}~x_{j_{2}}~\cdots x_{j_{2}+t-2}x_{j_{2}+t-1}\cdots x_{i_{2}-1}$
$\overline{x_{i_{2}+1}\cdots x_{i_{2}+t-1}}x_{i_{2}+t}\cdots x_{n}$
$=$ $x_{1}\cdots x_{i_{1}-1}~~x_{i_{1}}~\cdots x_{i_{1}+t-2}x_{i_{1}+t-1}\cdots x_{j_{1}-1}\overline{x_{j_{1}+1}\cdots}$
$\overline{x_{j_{1}+t-1}}x_{j_{1}+t}\cdots x_{j_{2}-1}\overline{x_{j_{2}+1}\cdots x_{j_{2}+t-1}}~~x_{j_{2}+t}\cdots~~x_{i_{2}}$~
$x_{i_{2}+1}\cdots x_{i_{2}+t-1}x_{i_{2}+t}\cdots x_{n}.$
Viewing $\mathbf{x}^{\prime}$, it is clear that two bursts of (1,t1)(1,t-1)-DS in $\mathbf{x}$ degrades into two deletions in a similar argument with Case i). Then, let $\mathcal{I}_{1}=[i_{1},j_{1}]$ and $\mathcal{I}_{2}=[j_{2},i_{2}]$, we have $|\mathcal{I}_{1}|,|\mathcal{I}_{2}|\leq d\log n$, then conclusion a) holds.
Case iii) $i_{1}+t\leq i_{2}\leq j_{1}-t\leq j_{2}-2t$. In this case,
$\mathbf{x}^{\prime}=$ $x_{1}\cdots x_{i_{1}-1}\overline{x_{i_{1}+1}\cdots x_{i_{1}+t-1}}~x_{i_{1}+t}~~\cdots x_{i_{2}-1}\overline{x_{i_{2}+1}\cdots}$
$\overline{x_{i_{2}+t-1}}~x_{i_{2}+t}~\cdots x_{j_{1}+1}x_{j_{1}+2}\cdots~~x_{j_{1}+t}~x_{j_{1}+t+1}\cdots$~~
$x_{j_{2}}~x_{j_{2}+1}\cdots x_{j_{2}+t-1}x_{j_{2}+t}\cdots x_{n}$
$=$ $x_{1}\cdots x_{i_{1}-1}~~x_{i_{1}}~\cdots x_{i_{1}+t-2}x_{i_{1}+t-1}\cdots x_{i_{2}-2}x_{i_{2}-1}\cdots$
$x_{i_{2}+t-3}x_{i_{2}+t-2}\cdots x_{j_{1}-1}\overline{x_{j_{1}+1}\cdots x_{j_{1}+t-1}}~~x_{j_{1}+t}$$\cdots$
$x_{j_{2}-1}\overline{x_{j_{2}+1}\cdots x_{j_{2}+t-1}}x_{j_{2}+t}\cdots x_{n}.$
That is, two bursts of (1,t1)(1,t-1)-DS in $\mathbf{x}$ degrades into the following form,
$\mathbf{x}^{\prime}=$ $x_{1}\cdots x_{i_{1}-1}x_{i_{1}+1}\cdots x_{i_{2}-1}\overline{x_{i_{2}+1}\cdots x_{i_{2}+t-1}}~~x_{i_{2}+t}$$\cdots$
$x_{j_{1}+1}x_{j_{1}+2}\cdots~~x_{j_{1}+t}~x_{j_{1}+t+1}\cdots~~x_{j_{2}}~x_{j_{2}+1}\cdots x_{n}$
$=$ $x_{1}\cdots x_{i_{1}-1}~~x_{i_{1}}~\cdots x_{i_{2}-2}x_{i_{2}-1}\cdots x_{i_{2}+t-3}x_{i_{2}+t-2}\cdots$
$x_{j_{1}-1}\overline{x_{j_{1}+1}\cdots x_{j_{1}+t-1}}~~x_{j_{1}+t}~\cdots x_{j_{2}-1}x_{j_{2}+1}\cdots x_{n}.$
By comparing the symbols of the corresponding positions, we obtain
$x_{i_{1}}=x_{i_{1}+1}=\cdots=x_{i_{2}-2}=x_{i_{2}-1},$ (30)
$x_{j_{1}+t}=x_{j_{1}+t+1}=\cdots=x_{j_{2}-1}=x_{j_{2}}.$
Meanwhile, if $j_{1}-i_{2}-t+1$ is even, then we have
$x_{i_{2}+t-2}=x_{i_{2}+t}=\cdots=x_{j_{1}-1}=x_{j_{1}+1},$ (31)
$x_{i_{2}+t-1}=x_{i_{2}+t+1}=\cdots=x_{j_{1}-2}=x_{j_{1}},$
if $j_{1}-i_{2}-t+1$ is odd, then
$x_{i_{2}+t-2}$ $=x_{i_{2}+t}=\cdots=x_{j_{1}-2}=x_{j_{1}},$ (32)
$x_{i_{2}+t-1}$ $=x_{i_{2}+t+1}=\cdots=x_{j_{1}-1}=x_{j_{1}+1}.$
Whether $j_{1}-i_{2}-t+1$ is even or odd, $\mathbf{x}_{[i_{2}+t-2,j_{1}+1]}$ is a run of $\mathbf{x}$ if $x_{j_{1}}=x_{j_{1}+1}$, otherwise it is an alternating substring in $\mathbf{x}$. In addition, $\mathbf{x}_{[i_{1},i_{2}-1]}$ and $\mathbf{x}_{[j_{1}+t,j_{2}]}$ are two runs of $\mathbf{x}$. Denoting $\mathcal{I}=[i_{1},j_{2}]$, we have $|\mathcal{I}|\leq 3d\log n+2t$ due to Equations (30)-(32) and Observation 1. Consequently, conclusion b) holds.
Furthermore, for scenarios involving overlapping bursts, the underlying intuition is illustrated in Figure 2. Following the same approach with the previous case i), two bursts of (1,t1)(1,t-1)-DS of $\mathbf{x}$ in Figure 2-i) and ii) degrades into two deletions, which can be handled correspondingly. Regarding to Figure 2-iii), we can observe that $\mathbf{x}_{[i_{1},i_{2}-1]}$ and $\mathbf{x}_{[j_{1}+t,j_{2}]}$ form two runs of $\mathbf{x}$. Hence, denoting $\mathcal{I}=[i_{1},j_{2}]$, conclusion b) follows. Since the details are close to those of the non-intersection cases, they are omitted for brevity.
Figure 2: Three different cases with intersections.
This completes the proof. ∎
Figure 3: The intervals $\mathcal{L}_{j},j\in[\lceil n/\rho\rceil-1]$ in (35)
The following conclusion is immediate from Lemma 6.
Lemma 7.
For a given binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$ and two positive integers $t_{1},t_{2}$ with $t_{1}>t_{2}$, let $X_{t_{1}-t_{2}}$ be the matrix representation of $\mathbf{x}$ following (25) and $f(\mathbf{x})$ be the function proposed in Lemma 3. Denote $t^{\prime}=\lceil t_{1}/(t_{1}-t_{2})\rceil$. Suppose that $X[1]$ is a $d$-regular sequence, then if $\mathbf{x}^{\prime}$ is obtained from $\mathbf{x}$ by two bursts of $(t_{1},t_{2})$-DI, one of the following conclusions hold.
  • a)
    One can find two intervals $\mathcal{J}_{1},\mathcal{J}_{2}\subseteq[n]$ with length $|\mathcal{J}_{1}|,|\mathcal{J}_{2}|\leq(d\log(n/(t_{1}-t_{2}))+2t^{\prime})(t_{1}-t_{2})$ such that one burst of $(t_{1},t_{2})$-DI occurs in $\mathbf{x}_{\mathcal{J}_{1}}$ and the other in $\mathbf{x}_{\mathcal{J}_{2}}$, respectively.
  • b)
    One can find an interval $\mathcal{J}\subseteq[n]$ with length $|\mathcal{J}|\leq(3d\log{(n/(t_{1}-t_{2}))}+4t^{\prime})(t_{1}-t_{2})$ such that two bursts of $(t_{1},t_{2})$-DI occur in $\mathbf{x}_{\mathcal{J}}$.
Proof.
Following Observation 2, two bursts of $(t_{1},t_{2})$-DI in $\mathbf{x}$ deduce two bursts of $(1,t^{\prime}-1)$-DS in $X[1]$. Denote $X^{\prime}_{t_{1}-t_{2}}$ as the matrix representation of $\mathbf{x}^{\prime}$. That is, it holds that $\mathbf{x}^{\prime}\in\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})$ and $X^{\prime}[1]\in\mathcal{B}_{2,(1,t^{\prime}-1)}^{DS}(X[1])$. According to Lemma 3, $X[1]$ can be recovered by $f(X[1])$ and $X^{\prime}[1]$. Moreover, since $X[1]$ is a $d$-regular sequence of length $n/(t_{1}-t_{2})$, we distinguish two cases following Lemma 6.
Case i) Lemma 6-a) holds. Denoting
$\mathcal{I}_{1}=[i,i^{\prime}],\mathcal{I}_{2}=[j,j^{\prime}]\subseteq\left[{n\over t_{1}-t_{2}}\right]$
with
$|\mathcal{I}_{1}|,|\mathcal{I}_{2}|\leq d\log{n\over t_{1}-t_{2}},$
by Observation 2, we know two bursts of $(t_{1},t_{2})$-DI occur in $\mathbf{x}_{\mathcal{J}_{1}}$ and $\mathbf{x}_{\mathcal{J}_{2}}$ with
$\mathcal{J}_{1}=\left[(i-t^{\prime})(t_{1}-t_{2})+2,(i^{\prime}+t^{\prime}-1)(t_{1}-t_{2})\right],$
$\mathcal{J}_{2}=\left[(j-t^{\prime})(t_{1}-t_{2})+2,(j^{\prime}+t^{\prime}-1)(t_{1}-t_{2})\right].$
It is clear that
$|\mathcal{J}_{1}|,|\mathcal{J}_{2}|\leq(d\log{n\over t_{1}-t_{2}}+2t^{\prime})(t_{1}-t_{2}).$
Case ii) Lemma 6-b) holds. Denoting
$\mathcal{I}=[i,i^{\prime}]\subseteq\left[{n\over t_{1}-t_{2}}\right],|\mathcal{I}|\leq 3d\log{n\over t_{1}-t_{2}}+2t^{\prime},$
we know two bursts of $(t_{1},t_{2})$-DI occur in $\mathbf{x}_{\mathcal{J}}$ with
$\mathcal{J}=\left[(i-t^{\prime})(t_{1}-t_{2})+2,(i^{\prime}+t^{\prime}-1)(t_{1}-t_{2})\right].$
It is clear that
$|\mathcal{J}|\leq\left(3d\log{n\over t_{1}-t_{2}}+4t^{\prime}\right)(t_{1}-t_{2}).$
Hence, the proof is completed. ∎
After approximately locating the positions of two bursts of $(t_{1},t_{2})$-DI, in what follows, we present an approach to recover the original sequence with the help of Theorem 5.
Let $\rho$ be a positive integer, and $\mathcal{L}_{j},j\in[\lceil n/\rho\rceil-1],$ be intervals of length $2\rho$, where
$\hskip-6.0pt\mathcal{L}_{j}=\left\{\begin{array}[]{ll}[(j-1)\rho+1,(j+1)\rho],&\hskip-8.0pt\text{for~~}j\in[\lceil n/\rho\rceil-2],\\ {[(j-1)\rho+1,n]},&\hskip-8.0pt\text{for~~}j=\lceil n/\rho\rceil-1.\end{array}\right$. (35)
We illustrate the partition of (35) in Figure 3.
Clearly, for any length-$\rho$ interval $\mathcal{L}\subseteq[n]$, there exists an $\mathcal{L}_{j_{1}},j_{1}\in[\lceil n/\rho\rceil-1],$ such that $\mathcal{L}\subseteq\mathcal{L}_{j_{1}}$.
Observation 3.
As shown in Figure 3, if there is one burst of $(t_{1},t_{2})$-DI in an interval $\mathcal{L}_{j},j\in[2,\lceil n/\rho\rceil-2]$, the adjacent intervals $\mathcal{L}_{j-1}$ and $\mathcal{L}_{j+1}$ might also be affected. Particularly, for $j=1$ and $j=\lceil n/\rho\rceil-1$, only $\mathcal{L}_{2}$ and $\mathcal{L}_{\lceil n/\rho\rceil-2}$ may be affected, respectively.
As a result, we provide a construction to correct two bursts of $(t_{1},t_{2})$-DI for binary sequences.
Construction 2.
For any binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$ and two positive integers $t_{1},t_{2}$ with $t_{1}>t_{2}$, denote by $X_{t_{1}-t_{2}}$ the matrix representation of $\mathbf{x}$ following (25), and $X[1]$ the first row of $X_{t_{1}-t_{2}}$. Let $f_{1}(\mathbf{x}),f_{2}(\mathbf{x})$ and $f(\mathbf{x})$ be the functions capable of correcting one burst of $(t_{1},t_{2})$-DI, two bursts of $(t_{1},t_{2})$-DI, and two bursts of $(1,t^{\prime}-1)$-DS proposed in Construction 1 and Lemma 3 for any $\mathbf{x}\in\Sigma_{2}^{n}$, respectively. Let $\mathcal{L}_{j}$ and $\mathcal{L}_{j^{\prime}}$ be the intervals defined in (35) with $\rho_{1}=(d\log(n/(t_{1}-t_{2}))+2t^{\prime})(t_{1}-t_{2})$ and $\rho_{2}=(3d\log(n/(t_{1}-t_{2}))+4t^{\prime})(t_{1}-t_{2})$, respectively, where $j\in[\lceil n/\rho_{1}\rceil-1/],j^{\prime}\in[\lceil n/\rho_{2}\rceil-1]$. For each $\mathbf{x}\in\Sigma_{2}^{n}$ with $X[1]$ being a $d$-regular sequence, $a\in[0,2]$ and $i\in[0,1]$, denote
$\phi^{(a)}_{i}(\mathbf{x})=\sum_{j\in[\lceil n/\rho_{1}\rceil-1],\atop j\equiv a\bmod 3}j^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j}})\pmod{2n^{i}N_{1}}$, (36)
$h^{(i)}(\mathbf{x})=\sum_{j\in[\lceil n/\rho_{1}\rceil-1],\atop j\equiv i\bmod 2}f_{2}(\mathbf{x}_{\mathcal{L}_{j}})\pmod{N_{2}}$, (37)
and
$g^{(i)}(\mathbf{x})=\sum_{j^{\prime}\in[\lceil n/\rho_{2}\rceil-1],\atop j^{\prime}\equiv i\bmod 2}f_{2}(\mathbf{x}_{\mathcal{L}_{j^{\prime}}})\pmod{N_{3}}$, (38)
where
$N_{1}=2^{4\log(2\rho_{1})+o(\log\rho_{1})},$
$N_{2}=2^{8\log(2\rho_{1})+o(\log\rho_{1})},$
and
$N_{3}=2^{8\log(2\rho_{2})+o(\log\rho_{2})}.$
Then, denote
$\psi(\mathbf{x})=\Big(f(X[1]),\phi^{(a)}_{i}(\mathbf{x}),h^{(i)}(\mathbf{x}),g^{(i)}(\mathbf{x}),$
$i\in[0,1],a\in[0,2]\Big).$
For some fixed parameters
$f\in\big[0,2^{8\log(n/(t_{1}-t_{2}))+o(\log n)}\big),$
$\phi_{i,a}\in[0,2n^{i}N_{1})$, $h_{i}\in[0,N_{2})$, and $g_{i}\in[0,N_{3})$, $a\in[0,2]$, $i\in[0,1]$, respectively, define a code by
$\mathcal{C}_{2}\triangleq\Big\{$ $\mathbf{x}\in\Sigma_{2}^{n}:\text{ $X[1]$ is a d-redular sequence},$
$\psi(\mathbf{x})=(f,\phi_{i,a},h_{i},g_{i}),i\in[0,1],a\in[0,2]\Big\}.$
Theorem 7.
The code $\mathcal{C}_{2}$ generated in Construction 2 can correct two bursts of $(t_{1},t_{2})$-DI with at most $11\log n+o(\log n)$ bits of redundancy.
Proof.
Based on the function $f(X[1])$ and $X^{\prime}[1]\in\mathcal{B}_{2,(1,t^{\prime}-1)}^{DS}(X[1])$, $X[1]$ can be recovered and the approximate locations of two bursts of $(t_{1},t_{2})$-DI can be identified following Lemmas 3 and 7. Then, we prove this result by considering the following two cases.
Case 1. Lemma 7-a) holds. In this case, we can find two intervals $\mathcal{J}_{1}$ and $\mathcal{J}_{2}$ with $|\mathcal{J}_{1}|,|\mathcal{J}_{2}|\leq\rho_{1}=(d\log(n/(t_{1}-t_{2}))+2t^{\prime})(t_{1}-t_{2})$ such that two bursts occur in $\mathbf{x}_{\mathcal{J}_{1}}$ and $\mathbf{x}_{\mathcal{J}_{2}}$, respectively. Following (35), there exist two intervals $\mathcal{L}_{j_{1}},\mathcal{L}_{j_{2}}$ such that $\mathcal{J}_{1}\subseteq\mathcal{L}_{j_{1}},\mathcal{J}_{2}\subseteq\mathcal{L}_{j_{2}}$, where $j_{1},j_{2}\in[\lceil n/\rho_{1}\rceil-1]$. Without loss of generality, let $j_{1}\leq j_{2}$ in the following.
We consider the following two subcases:
1) $j_{1}\neq j_{2}\in[2,\lceil n/\rho_{1}\rceil-2]$. Assume that $\mathcal{L}_{j_{1}}=[\lambda_{0},\lambda_{1}]$, $\mathcal{L}_{j_{2}}=[\lambda_{2},\lambda_{3}]$. Thus, we have
$\mathbf{x}_{[1,\lambda_{0})}=\mathbf{x}^{\prime}_{[1,\lambda_{0})},$
$\mathbf{x}_{[\lambda_{1}+1,\lambda_{2})}=\mathbf{x}^{\prime}_{[\lambda_{1}-t_{1}+t_{2}+1,\lambda_{2}-t_{1}+t_{2})},$ (39)
$\mathbf{x}_{[\lambda_{3}+1,n]}=\mathbf{x}^{\prime}_{[\lambda_{3}-2t_{1}+2t_{2}+1,n-2t_{1}+2t_{2}]},$
and
$\begin{array}[]{lll}\mathbf{x}^{\prime}_{[\lambda_{0},\lambda_{1}-t_{1}+t_{2}]}\in\mathcal{B}_{(t_{1},t_{2})}^{DI}(\mathbf{x}_{\mathcal{L}_{j_{1}}}),\\ \mathbf{x}^{\prime}_{[\lambda_{2}-t_{1}+t_{2},\lambda_{3}-2t_{1}+2t_{2}]}\in\mathcal{B}_{(t_{1},t_{2})}^{DI}(\mathbf{x}_{\mathcal{L}_{j_{2}}}).\end{array}$ (42)
Denoting $j_{1}\bmod 3=m_{1}$ and $j_{2}\bmod 3=m_{2}$, according to Observation 3 and (36), for any $a\in[0,2],i\in[0,1]$, we can obtain
i) when $a=m_{1}=m_{2}$,
$j_{1}^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j_{1}}})+j_{2}^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j_{2}}})$
$=$ $\phi^{(a)}_{i}(\mathbf{x})-\sum_{j\equiv a\bmod 3,\atop j\notin\{j_{1},j_{2}\}}j^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j}})\pmod{2n^{i}N_{1}}$. (43)
Note that since it is evident that $j\in[\lceil n/\rho_{1}\rceil-1]$, we omit this condition hereafter for notational simplicity.
ii) when $a=m_{1}\equiv m_{2}-1\pmod{3}$,
$j_{1}^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j_{1}}})+(j_{2}-1)^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j_{2}-1}})$
$=$ $\phi^{(a)}_{i}(\mathbf{x})-\sum_{j\equiv a\bmod 3,\atop j\notin\{j_{1},j_{2}-1\}}j^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j}})\pmod{2n^{i}N_{1}}$, (44)
and
$(j_{1}+1)^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j_{1}+1}})+j_{2}^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j_{2}}})$
$=$ $\phi^{(a+1)}_{i}(\mathbf{x})-\sum_{j\equiv a\bmod 3,\atop j\notin\{j_{1}+1,j_{2}\}}j^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j}})\pmod{2n^{i}N_{1}}$. (45)
Note that the superscript $a+1$ of $\phi_{i}$ needs to module $3$.
iii) when $a=m_{1}\equiv m_{2}-2\pmod{3}$,
$j_{1}^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j_{1}}})+(j_{2}+1)^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j_{2}+1}})$
$=$ $\phi^{(a)}_{i}(\mathbf{x})-\sum_{j\equiv a\bmod 3,\atop j\notin\{j_{1},j_{2}+1\}}j^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j}})\pmod{2n^{i}N_{1}}$, (46)
and
$(j_{1}-1)^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j_{1}-1}})+j_{2}^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j_{2}}})$
$=$ $\phi^{(a+2)}_{i}(\mathbf{x})-\sum_{j\equiv a\bmod 3,\atop j\notin\{j_{1}-1,j_{2}\}}j^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j}})\pmod{2n^{i}N_{1}}$. (47)
Note that the superscript $a+2$ of $\phi_{i}$ needs to module $3$.
From Theorem 5, we know $f_{1}(\mathbf{x}_{\mathcal{L}_{j}})\leq N_{1},j\in[\lceil n/\rho_{1}\rceil-1]$. Therefore, regardless of which of the three subcases above applies, $f_{1}(\mathbf{x}_{\mathcal{L}_{j_{1}}})$ and $f_{1}(\mathbf{x}_{\mathcal{L}_{j_{2}}})$ can be solved based on (V-B)-(V-B). Since two bursts of $(t_{1},t_{2})$-DI are in $\mathbf{x}_{\mathcal{L}_{j_{1}}}$ and $\mathbf{x}_{\mathcal{L}_{j_{2}}}$, the original sequence $\mathbf{x}$ can be uniquely recovered following (V-B), (42), and Theorem 5.
Particularly, for the case of $j_{1}=1$, the Equation (V-B) shows
$\begin{aligned} j_{2}^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j_{2}}})=\phi^{(a+2)}_{i}(\mathbf{x})-\sum_{j\equiv a\bmod 3,\atop j\neq j_{2}}j^{i}\cdot f_{1}(\mathbf{x}_{\mathcal{L}_{j}})\pmod{2n^{i}N_{1}}\end{aligned}$
and for the case of $j_{2}=\lceil n/\rho_{1}\rceil-1$, the Equation (V-B) gives
[…]
This means that there are also at most two unknowns in equations (V-B)-(V-B), which can be solved correspondingly.
2) $j_{1}=j_{2}\in[\lceil n/\rho_{1}\rceil-1]$. In this subcase, two bursts of $(t_{1},t_{2})$-DI occur in a length-$2\rho_{1}$ substring $\mathbf{x}_{\mathcal{L}_{j_{1}}}$ of $\mathbf{x}$. Assume that $\mathcal{L}_{j_{1}}=[\lambda_{0},\lambda_{1}]$, thus we know
$\begin{array}[]{lll}\mathbf{x}_{[1,\lambda_{0})}=\mathbf{x}^{\prime}_{[1,\lambda_{0})},\\ \mathbf{x}_{[\lambda_{1}+1,n]}=\mathbf{x}^{\prime}_{[\lambda_{1}-2t_{1}+2t_{2}+1,n-2t_{1}+2t_{2}]},\end{array}$ (50)
and
$\mathbf{x}^{\prime}_{[\lambda_{0},\lambda_{1}-2t_{1}+2t_{2}]}\in\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x}_{\mathcal{L}_{j_{1}}}).$ (51)
According to (37), while $j_{1}\bmod 2=0$, we can compute
$f_{2}(\mathbf{x}_{\mathcal{L}_{j_{1}}})=h^{(0)}(\mathbf{x})-\sum_{j\bmod 2=0,\atop j\neq j_{1}}f_{2}(\mathbf{x}_{\mathcal{L}_{j}})\pmod{N_{2}}$, (52)
while $j_{1}\bmod 2=1$,
$f_{2}(\mathbf{x}_{\mathcal{L}_{j_{1}}})=h^{(1)}(\mathbf{x})-\sum_{j\bmod 2=1,\atop j\neq j_{1}}f_{2}(\mathbf{x}_{\mathcal{L}_{j}})\pmod{N_{2}}$. (53)
From Theorem 5, we know $f_{2}(\mathbf{x}_{\mathcal{L}_{j}})\leq N_{2},j\in[\lceil n/\rho_{1}\rceil-1]$. Since $\mathbf{x^{\prime}}\in\mathcal{B}_{2,(t_{1},t_{2})}^{DI}(\mathbf{x})$, by means of $f_{2}(\mathbf{x}_{\mathcal{L}_{j_{1}}})$ in (52) or (53), the original sequence $\mathbf{x}$ can be uniquely recovered following (50), (51), and Theorem 5.
Case 2. Lemma 7-b) holds. In this case, we can find an interval $\mathcal{J}\subseteq[n]$ with length $|\mathcal{J}|\leq\rho_{2}=(3d\log(n/(t_{1}-t_{2}))+4t^{\prime})(t_{1}-t_{2})$ such that two bursts of $(t_{1},t_{2})$-DI occur in $\mathbf{x}_{\mathcal{J}}$. Following (35), we know that there exist an interval $\mathcal{L}_{j}$ such that $\mathcal{J}\subseteq\mathcal{L}_{j}$, where $j\in[\lceil n/\rho_{2}\rceil-1]$. Then, the proof is similar to that of subcase 2) in Case 1, with the only difference being the use of $g^{(i)}(\mathbf{x})$ in place of $h^{(i)}(\mathbf{x})$. The details are therefore omitted.
Following Lemma 3, the equations (36)-(38), the length of $\psi(\mathbf{x})$ satisfies
$\begin{aligned} \left|\psi(\mathbf{x})\right|=&\left|f(X[1])\right|+\sum_{i=0}^{1}\sum_{a=0}^{2}\left|\phi^{(a)}_{i}(\mathbf{x})\right|+\sum_{i=0}^{1}\left|h^{(i)}(\mathbf{x})\right|\\ &+\sum_{i=0}^{1}\left|g^{(i)}(\mathbf{x})\right|\\ =&8\log{n\over t_{1}-t_{2}}+o(\log n)+3\left(\log(2N_{1})+\log(2nN_{1})\right)\\ &+2(\log N_{2})+2(\log N_{3})\\ =&11\log{n}+o(\log n).\end{aligned}$
completing this proof. ∎
According to Lemma 5, and Theorems 6 and 7, the following conclusion is straightforward.
Theorem 8.
For any binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$ and two positive integers $t_{1},t_{2}$ with $t_{1}\geq t_{2}$, two bursts of $(t_{1},t_{2})$-DI in $\mathbf{x}$ can be corrected with at most $11\log n+o(\log n)$ bits of redundancy.
V-C Complexity
In Construction 1, two bursts of $(t_{1},t_{2})$-DI errors in a binary sequence $\mathbf{x}$ are corrected by the syndrome compression technique directly. Regarding to Construction 2, we firstly locate two bursts of $(1,t^{\prime}-1)$-DS in the first row of matrix representation of $\mathbf{x}$ using a function achieved by the syndrome compression technique, and then correct two bursts of $(t_{1},t_{2})$-DI of $\mathbf{x}$. As shown in [24], the main computational cost of the syndrome compression technique arises from a brute-force search for a parameter
$a(\mathbf{x})\leq 2^{\log{\left|\mathcal{N}^{*}(\mathbf{x})\right|+o(\log n)}}$
such that
$a(\mathbf{x})\notin\big\{j:j\mid\big(h(\mathbf{x})-h(\mathbf{x}^{\prime})\big)\big\}$
for any $\mathbf{x}^{\prime}\in\mathcal{N}^{*}(\mathbf{x})$, where $\mathcal{N}^{*}\in\{\mathcal{N}^{DI}_{2,(t_{1},t_{2})},\mathcal{N}^{DS}_{2,(1,t^{\prime}-1)}\}$ and $h(\mathbf{x})$ is defined in Lemma 2. Therefore, the computational complexity of this technique is directly determined by the size of $\mathcal{N}^{*}(\mathbf{x})$. In what follows, we compute the cardinality of these two sets. Specifically, since the length of $X[1]$ is $n/(t_{1}-t_{2})$, from (23), the cardinality of $\mathcal{N}^{DS}_{2,(1,t^{\prime}-1)}(\mathbf{x})$ in Lemma 3 is
$\begin{aligned} \left|\mathcal{N}^{DS}_{2,(1,t^{\prime}-1)}(\mathbf{x})\right|=&{{n\over t_{1}-t_{2}}-2t^{\prime}+2\choose 2}{{n\over t_{1}-t_{2}}-2t^{\prime}+3\choose 2}\Big(\sum_{i=0}^{t^{\prime}-2}{t^{\prime}-2\choose i}\Big)^{4}\\ \leq&O\left(\left({n\over{t_{1}-t_{2}}}\right)^{4}\right),\end{aligned}$
while the one in Construction 1 is
$\begin{aligned} \left|\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})\right|=&{n-2t_{1}+2\choose 2}{n-2t_{1}+2t_{2}+1\choose 2}\cdot 2^{2t_{1}+2t_{2}-8}\\ \leq&O(n^{4})\end{aligned}$
for $t_{1},t_{2}\geq 2$, and
$\begin{aligned} \left|\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})\right|=&{n-2t_{1}+2\choose 2}{n-2t_{1}+2t_{2}+1\choose 2}\cdot 2^{2t_{1}-4}\\ \leq&O(n^{4})\end{aligned}$
for $t_{1}>t_{2}=1$.
Moreover, for the case of $t_{1}>t_{2}$, the functions $\phi^{(a)}_{i}(\mathbf{x})$, $h^{(i)}(\mathbf{x})$ and $g^{(i)}(\mathbf{x})$ in Construction 2 are computable in time $O((2\rho_{1})^{4})$, $O((2\rho_{1})^{2})$ and $O((2\rho_{2})^{4})$, where $\rho_{1}=(d\log(n/(t_{1}-t_{2}))+2t^{\prime})(t_{1}-t_{2})$, and $\rho_{2}=(3d\log(n/(t_{1}-t_{2}))+4t^{\prime})(t_{1}-t_{2})$, respectively. This means that the functions $\phi^{(a)}_{i}(\mathbf{x}),h^{(i)}(\mathbf{x})$ and $g^{(i)}(\mathbf{x})$ have at most $O({(\log n)}^{4})$ time complexity. Hence, the computational complexity of Construction 2 is mainly from $f(X[1])$, i.e., the cardinality of $\mathcal{N}^{DS}_{2,(1,t^{\prime}-1)}(\mathbf{x})$. Consequently, comparing (54) to (55) and (56), the Construction 2 reduces computational complexity by a factor of $(t_{1}-t_{2})^{4}$.
In addition, for the case of $t_{1}=t_{2}$, two bursts of $(t_{1},t_{2})$-DI errors are transferred to four substitutions, which are corrected by a GRS code. Both of the encoding and decoding complexities for GRS codes are $O(n\log n)$, which is extremely less than $O(n^{4})$.
For clarity, we compare the cardinality of $\mathcal{N}^{*}(\mathbf{x})$ for Constructions 1 and 2 by a specific example. Note that the numerical values below are not exact, since they are evaluated for finite $n$, while the complexity analysis is asymptotic in nature (i.e., assumes $n\rightarrow\infty$). Nevertheless, we provide them to elucidate computational complexity benefits of the new constructions. Given $n=256,t_{1}=10,t_{2}=2$, then $t^{\prime}=2$. For any binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$, if two bursts of $(t_{1},t_{2})$-DI occur in $\mathbf{x}$, the cardinality of $\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})$ in (55) is
$|\mathcal{N}_{2,(10,2)}(\mathbf{x})|={238\choose 2}{241\choose 2}\times 2^{16}\approx 2^{46},$ (57)
while the one in (54) is
$|\mathcal{N}^{DS}_{2,(1,1)}(\mathbf{x})|={30\choose 2}{31\choose 2}\approx 2^{20}.$ (58)
It is clear that the computational complexity associated with (58) is significantly lower than that of (57), indicating that the practical applicability of Construction 1 is limited due to the substantially higher computational cost. In contrast, Construction 2 exhibits superior computational efficiency and is therefore more amenable to implementation.
TABLE I: Comparisons of the codes in Thm. 5 and Thm. 8 under some fixed parameters
Row Codelength $n$ $(t_{1},t_{2})$-DI $t^{\prime}$ $\left|\mathcal{N}^{*}(\mathbf{x})\right|$
Thm. 5 Thm. 8
1 $1024$ $(10,8)$ 55 $2^{66}$ $2^{46}$
2 $512$ $2^{62}$ $2^{42}$
3 $256$ $2^{58}$ $2^{38}$
4 $1024$ $(10,7)$ 44 $2^{64}$ $2^{40}$
5 $512$ $2^{60}$ $2^{36}$
6 $256$ $2^{56}$ $2^{32}$
7 $1024$ $(10,6)$ 33 $2^{62}$ $2^{34}$
8 $512$ $2^{58}$ $2^{30}$
9 $256$ $^{54} 22$$^{26}$
10 $1024$ $(10,1)$ 22 $^{54} 22$$^{26}$
11 $512$ $^{50} 22$$^{22}$
12 $256$ $2^{46}$ $2^{18}$
13 $\left|\mathcal{N}_{2,(t_{1},t_{2})}(\mathbf{x})\right|$ Operations of GRS codes
14 $512$ $(10,10)$ - $2^{74}$ $2^{15}$
15 $512$ $(5,5)$ $2^{54}$ $2^{17}$
16 $512$ $(2,2)$ $2^{34}$ $2^{19}$
Furthermore, a numerical comparison between the codes in Theorems 5 and 8 for several fixed parameters is presented in Table I. It can be observed that, across all parameter settings, the computational complexity of those in Theorem 8 is extremely lower than the one in Theorem 5. For instance, when $n=1024,512,256$, if there are two bursts of $(t_{1}=10,t_{2}=6)$-DI occur in a binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$, the codes in Theorem 8 yield a significant reduction in the size of $\mathcal{N}^{DS}_{2,(1,2)}(\mathbf{x})$, improving from $\left|\mathcal{N}^{DI}_{2,(10,6)}(\mathbf{x})\right|\approx 2^{62},2^{58}$ and $2^{54}$ in Theorem 5 to $\left|\mathcal{N}^{DS}_{2,(1,2)}(\mathbf{x})\right|\approx 2^{34},2^{30}$ and $2^{26}$, respectively, a constant factor improvement of $2^{28}$ across all code lengths (Rows 7-9).
VI Conclusions
In this paper, we developed binary ECCs capable of correcting two bursts of $(t_{1},t_{2})$-DI errors. We first established that two bursts of $(t_{1},t_{2})$-DI ECCs, two bursts of $(t_{2},t_{1})$-DI ECCs, and one burst of $(t_{1},t_{2})$-DI together with one burst of $(t_{2},t_{1})$-DI ECCs are equivalent. Leveraging this equivalence, we characterized lower and upper bounds on the code size of two bursts of $(t_{1},t_{2})$-DI ECCs, thereby showing that the redundancy of such codes satisfies $2\log n+O(1)\leq r(\mathcal{C})\leq 4\log n+o(\log n)$. Furthermore, focusing on the cases with $t_{1}=t_{2}$ and $t_{1}>t_{2}$, we presented constructions of binary two bursts of $(t_{1},t_{2})$-DI ECCs with redundancy at most $8\log n+o(\log n)$ and $11\log n+o(\log n)$, respectively. The latter was achieved by transforming two bursts of $(t_{1},t_{2})$-DI errors in a binary sequence into two bursts of $(1,t^{\prime}-1)$-DS errors in the first row of its matrix representation, where $t^{\prime}=\lceil t_{1}/(t_{1}-t_{2})\rceil$, coupled with the syndrome compression technique. Compared with a direct application of syndrome compression, the proposed codes achieve a significant reduction in computational complexity.